### v1

# A Comprehensive Evaluation of Large Language Models on Mental Illnesses

Abstract
Large Language Model (LLM) s have shown promise in various domains, including healthcare. In this study, we conducted a comprehensive evaluation of LLM s in the context of mental health tasks using social media data. We explored the zero-shot (ZS) and few-shot (FS) capabilities of various LLM s, including GPT-4, Llama 3, Claude, Gemma, Gemini, Phi-3, and others, on tasks such as binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment. Our evaluation involved 33 models ranging from 2 billion to 405+ billion parameters, and we utilized 9 main prompt templates across the tasks. Key findings revealed that models like GPT-4 and Llama 3 exhibited superior performance in binary disorder detection, with accuracies reaching up to 85% on certain datasets. Moreover, prompt engineering played a crucial role in enhancing model performance. For example, when transitioning from our first binary prompt to a more structured prompt, the accuracy of certain models increased significantly. Notably, the Mixtral 8x22b model showed an improvement of over 20%, while Gemma 7b experienced a similar boost in performance. In the task of disorder severity evaluation, we observed that FS learning significantly improved the model’s accuracy, highlighting the importance of contextual examples in complex assessments. Notably, the Phi-3-mini model exhibited a substantial increase in performance, with balanced accuracy (BA) improving by over 6.80% and mean average error (MAE) dropping by nearly 1.3 when moving from ZS to FS learning. Additionally, in the psychiatric knowledge assessment task, recent models generally outperformed older, larger counterparts, with the Llama 3.1 405b achieving an accuracy of 91.2%. Despite promising results, our analysis identified several challenges, including variability in performance across datasets and the need for careful prompt engineering. There is also the high cost associated with using large models and the limitations imposed by the quality of social media data. Furthermore, the ethical guards imposed by many LLM providers hamper the ability to accurately evaluate their performance, due to tendency to not respond to potentially sensitive queries.
Keywords Large Language Models, Mental Health, Psychiatry, Social Media, Natural Language Processing, Zero-shot Learning, Few-shot Learning, Prompt Engineering, Model Evaluation
1 Introduction
Artificial Intelligence (AI) is rapidly transforming the landscape of mental healthcare, offering innovative solutions to address the growing global burden of mental illness. According to the World Health Organization, mental disorders accounted for 5.1% of the global disease burden in 2019, with depressive disorders affecting 280 million people worldwide, and an estimated 703,000 people dead by suicide [1]. The economic consequences of mental health conditions are also significant. In the United States, the cumulative cost of mental health inequities is projected to reach a staggering US\$14 trillion between 2024 and 2040. These costs include direct medical expenses, emergency department utilization, productivity losses, and premature deaths [2].
The integration of AI in psychiatry has enabled a wide array of applications, spanning from the crucial tasks of early detection and diagnosis to predicting treatment outcomes and providing therapeutic interventions. AI models have been employed to analyze diverse data modalities, including neuroimaging data such as functional magnetic resonance imaging (fMRI) and electroencephalogram (EEG), physiological signals like electrocardiogram (ECG), genetic information, and demographic data [3].
The advent of Natural Language Processing (NLP) in mental health research opened up new possibilities. Early NLP applications focused on analyzing text from clinical interviews and self-report questionnaires. These applications used techniques like the Linguistic Inquiry and Word Count to quantify linguistic patterns associated with various mental health conditions [4]. The emergence of transformer-based models, such as Bidirectional Encoder Representations from Transformers (BERT) [5], marked a significant turning point in NLP for mental health. These models, trained on massive amounts of text data, can automatically learn complex linguistic representations, eliminating the need for manual feature engineering. This has led to substantial improvements in tasks like sentiment analysis, suicide risk assessment, and the detection of mental health conditions from social media posts [6].
The advancement to LLM s marks a significant leap forward in the field of NLP. LLM s, such as OpenAI’s GPT-4 [7], Google’s Gemini [8, 9], Meta’s Llama [10], and many others have been trained on vast corpora of text data from diverse sources across the internet, encompassing billions of words and a wide range of human knowledge. This extensive training enables LLM s to understand and generate human-like text with a remarkable degree of fluency and coherence.
Given their extensive training, LLM s have immense potential in the realm of mental health and psychiatry. They can be employed for initial consultations, providing patients with preliminary information about mental health disorders and guiding them toward appropriate care pathways. LLM s can quickly analyze transcripts from patient sessions, offering initial insights and supporting medical practitioners in making more informed decisions. Their ability to process large volumes of patient history and previous medication data can streamline administrative tasks, ensuring that practitioners have quick access to comprehensive patient profiles. Additionally, LLM s can serve as virtual counselors, offering support and guidance to individuals experiencing poor mental states, thus extending the reach of mental health services to those who may not have immediate access to professional help. These models can also assist in monitoring social media for signs of mental distress, enabling timely interventions and potentially saving lives. As these technologies continue to evolve, their integration into mental healthcare promises to enhance the efficiency, accessibility, and quality of care provided to patients [11].
However, most if not all of these capabilities have yet to be properly verified for use in any real medical context. While the potential is significant, rigorous validation and clinical trials are necessary to ensure their effectiveness and safety in practical applications. Current work in the literature revolves around a few key topics, including evaluation (Sec. 1.1), fine-tuning (Sec. 1.2), data augmentation and chatbots (Sec. 1.3), benchmarks (Sec. 1.4), and literature reviews (Sec. 1.5).
1.1 Evaluating LLM s on Mental Health Tasks
One of the earliest works in this domain is the study by [12], which presents a simple evaluation of GPT-3.5-Turbo’s ZS binary classification performance on stress, depression, and suicidality detection tasks using social media posts. The resulting F1 scores (73% for stress, 86% for depression, and 37% for suicide), were particularly bad in the suicidality severity prediction task. The authors of [13] extend this evaluation by assessing GPT-3.5-Turbo’s text classification abilities on a broader range of affective computing problems, including big-five personality prediction and sentiment analysis, in addition to suicide tendency detection. Their findings reveal that while GPT-3.5-Turbo provided decent results, comparable to Word2Vec and bag-of-words baselines, it did not outperform the fine-tuned RoBERTa model on the given tasks. However, the authors did not prompt GPT-3.5-Turbo through an Application Programming Interface (API), they instead did it manually using the web interface, only resetting the context every 25 prompts, which might have affected their results.
1.2 Fine-tuning LLM s for mental health tasks
Building upon the evaluation studies, [14] and [15] explore fine-tuning LLM s for improved performance in mental health tasks. [14] introduce the MentaLlama model, which is fine-tuned on the IMHI dataset, a large dataset of social media posts labeled with mental health conditions and explanations. The authors created the IMHI dataset by taking multiple existing mental health datasets and generating additional explanatory labels using GPT-3.5-Turbo. The MentaLlama-chat-13B model surpasses or approaches state-of-the-art discriminative methods in prediction correctness on 7 out of 10 test sets and generates explanations on par with GPT-3.5-Turbo thus providing interpretability.
The authors of [15] conducted a comprehensive evaluation of multiple LLM s, including Alpaca, Alpaca-LoRA, FLAN-T5, Llama 2, GPT-3.5-Turbo, and GPT-4, on various mental health prediction tasks using online text data. Their work explored ZS prompting, FS prompting, and instruction fine-tuning techniques. The authors fine-tuned Alpaca and FLAN-T5 on six mental health prediction tasks across four datasets from Reddit. Results show that instruction fine-tuning significantly improves the performance of LLM s on these tasks. The fine-tuned models, Mental-Alpaca and Mental-FLAN-T5, outperformed much larger models like GPT-3.5-Turbo and GPT-4 on specific tasks, and performed on par with the state-of-the-art task-specific model Mental-RoBERTa [6]. The study also includes an exploratory case study on these models’ reasoning capability, which further suggests both the promising future and the important limitations of LLM s.
1.3 LLM s for Data Augmentation and Chatbot Development
[16], [17], and [18] explore the application of LLM s beyond prediction tasks, focusing on data augmentation and chatbot development. [16] develop Chat-Diagnose, an explainable and interactive LLM-augmented system for depression detection in social media. Chat-Diagnose incorporates a tweet selector to filter excessive posts, an image descriptor to convert images into text, and professional diagnostic criteria (Diagnostic and Statistical Manual of Mental Disorders (DSM)-5) to guide the diagnostic process. The system utilizes the chain-of-thought (CoT) technique to provide explanations and diagnostic evidence. In the full-training setting, it leverages answer heuristics from a traditional depression detection model. Chat-Diagnose achieves state-of-the-art (SoTA) performance in various settings, including ZS, FS, and full-training scenarios.
[17] investigates the use of GPT-3.5-Turbo for generating synthetic Reddit posts simulating depression symptoms from the Beck Depression Inventory (BDI)-II questionnaire, aiming to enhance semantic search capabilities. The study finds that using original BDI-II responses as queries is more effective, suggesting that the generated data might be too specific for this task.
[18] focuses on developing chatbots that simulate psychiatrists and patients in clinical diagnosis scenarios, specifically for depressive disorders. The study follows a three-phase design, involving collaboration with psychiatrists, experimental studies, and evaluations with real psychiatrists and patients. The development process emphasizes an iterative refinement of prompt design and evaluation metrics based on feedback from both psychiatrists and patients. A key finding of the study is the comparison between real and simulated psychiatrists, revealing differences in questioning strategies and empathy behaviors.
1.4 Benchmarks for Evaluating LLM s in Psychiatry
In contrast to the aforementioned studies, the authors of [19] took a different approach. Instead of evaluating existing LLM s on established datasets, they sought to create a comprehensive benchmark specifically designed to assess the capabilities of LLM s in the mental health domain. PsyEval (the benchmark) includes tasks such as mental health question-answering, where LLM s are evaluated on their ability to provide accurate and informative responses to queries related to mental health conditions, symptoms, and treatments. The benchmark also features tasks focused on diagnosis prediction, both using data from social media or simulated dialouges between patients and clinicians. Additionally, PsyEval incorporates tasks that assess LLM s’ ability to provide empathetic and safe psychological counseling in simulated conversations.
1.5 Literature Reviews
In addition to the contribution papers mentioned above, several surveys have been conducted to summarize and analyze the landscape of LLM s in mental health care. The authors of [11] provide a comprehensive overview of the opportunities and risks associated with LLM s in psychiatry, discussing their potential to enhance mental health care through improved diagnostic accuracy, personalized care, and streamlined administrative processes. The authors highlight the ability of LLM s to efficiently analyze patient data, summarize therapy sessions, and aid in complex diagnostic problem-solving. They also discuss the potential of LLM s to automate certain managerial decisions within hospital systems, such as personnel schedules and equipment needs. However, the authors also acknowledge the risks associated with LLM s, such as labor substitution, the potential for reduced human-to-human socialization, and the amplification of existing biases. They conclude by advocating for the development of pragmatic frameworks, including red-teaming and multi-stakeholder-oriented safety and ethical guidelines, to ensure the safe and responsible deployment of LLM s in psychiatry.
The authors of [20] conducted a systematic review of the applications of LLM s, such as GPT-3.5-Turbo, in the field of psychiatry. The review identified 16 studies that directly examined the use of LLM s in psychiatry, with applications ranging from clinical reasoning and social media analysis to education. The authors found that LLM s like GPT-3.5-Turbo and GPT-4 showed promise in assisting with tasks such as diagnosing mental health issues, managing depression, evaluating suicide risk, and supporting education in the field. However, the review also highlighted limitations, such as difficulties with complex cases and potential underestimation of suicide risks.
The article by [21] presents a scoping review of LLM s in mental health care, focusing on their applications and outcomes. The review identified 34 publications that met their inclusion criteria, including studies compiling new datasets, studies focused on model development, studies focused on employing and evaluating LLM s, and studies examining the potential opportunities, risks and ethical considerations associated with utilizing LLM s in a psychiatric context. The authors provide a detailed overview of the models and datasets used in these studies. They found that most studies utilized GPT-3.5 or GPT-4, with some fine-tuning on mental health-specific datasets like MentaLlama and Mental-LLM. The datasets used in these studies were primarily sourced from social media platforms and online mental health forums, with a few utilizing clinical data or simulated conversations. The authors also discuss the key challenges associated with LLM s, such as data availability and reliability, effective handling of mental states, and appropriate evaluation methods.
1.6 Research Gaps
While the potential of LLM s in mental health care is evident, several research gaps remain to be addressed. Firstly, most studies have employed relatively outdated models like GPT-3.5-Turbo, GPT-4, and Llama 2, with no investigations into newer models such as GPT-4-0, GPT-4-Turbo, Llama 3, Phi-3, Gemma 1 & 2, Claude models, and others. This limits our understanding of how the latest advancements in LLM technology can be leveraged for mental health applications. Secondly, there is a lack of focus on prompt engineering to investigate the effect of different prompts on disorder detection tasks. While a few studies have touched upon this aspect, a more systematic and comprehensive exploration of prompt engineering techniques is needed to optimize LLM performance in this domain. Thirdly, many studies have overlooked smaller yet important factors such as LLM variation across runs and the effect of small prompt modifications across models. Additionally, there is often a lack of full transparency in reporting methodologies, even when code is provided, hindering reproducibility and further research.
The goal of this work is to comprehensively evaluate the capabilities of LLM s on tasks relating to the detection of mental illness symptoms or mental illnesses from social media. We focus on several unique contributions:
  • We test models never used before in this context, including Gemini 1.0 Pro, GPT-4/GPT-4-Turbo/GPT-4o/GPT-4o mini, Mistral NeMo/Medium/Large/Large 2, Claude 3/3.5, Mixtral, Gemma 1/2, Phi-3, Llama 3/3.1, and Qwen 2.
  • We conduct an investigation of published human-annotated mental health datasets, both private and public.
  • We conduct extensive experiments with various prompting methods, including FS prompting and severity prompting.
  • We perform unique experiments such as temperature variation tests, filtering strategies, and chain-of-thought prompting to explore model performance and interpretability.
  • We provide comprehensive details about our methodology, including all prompts, to ensure full transparency and facilitate reproducibility.
  • We provide a detailed list of drawbacks, and challenges that currently exist when exploring the applications of LLM s in psychiatry, along with suggestions for future researchers.
The rest of this paper is organized as follows. Section 2 details our methodology implementation, including the experimental design, datasets, models, prompt templates, evaluation metrics, and parsing of model outputs. Section 3 presents and discusses the results of our experiments, analyzing the performance of various LLM s across the three tasks. Section 4 delves into additional investigations and experiments conducted alongside our main analysis. Finally, Section 5 concludes the paper by summarizing our findings, discussing the challenges encountered, and proposing potential directions for future research.
2 Methodology Implementation
To investigate the potential of LLM s in mental health tasks, we designed a series of experiments to evaluate their performance across various scenarios. This section details our experimental setup, including the chosen datasets, models, prompting strategies, and evaluation metrics.
2.1 Experimental Setup
Our experiments were designed to assess LLM capabilities across three primary tasks: binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment. We conducted ZS experiments for all three tasks, where the LLM s were given only the input data (social media post or multiple-choice question) and the task description, without any prior examples. For the disorder severity prediction task, we also conducted FS experiments, where the LLM s were provided with three examples of posts and their corresponding severity ratings before being asked to predict the severity of a new post.
2.1.1 Task 1: ZS Binary Disorder Detection
ZS learning is a machine learning paradigm where the model makes predictions without having seen any labeled examples of the specific task during training. Instead, the model relies on its pre-existing knowledge and the task description to make inferences. This approach is particularly useful when labeled data is scarce or when deploying models to new tasks quickly.
In this analysis, ZS learning was employed to evaluate the LLM’s inherent understanding of psychiatric disorders. Each LLM was tasked with determining whether a social media user exhibits a specific mental disorder based solely on their post and a brief task description. We focused on three prevalent disorders: depression, suicide risk, and stress. This setup allowed us to assess the model’s ability to identify these disorders based on its pre-trained knowledge and general language understanding.
2.1.2 Task 2: ZS/FS Disorder Severity evaluation
FS learning, on the other hand, is a technique where the model is provided with a small number of labeled examples before being tasked with making predictions. This approach helps the model to better understand the context and complexities of the task.
This task involved evaluating the severity (e.g. from 0 to 4) of a specific mental disorder in a user solely from a social media post written by them. We evaluated LLM s on depression and suicide risk, employing both ZS and FS approaches. In the ZS scenario, the LLM received only the social media post and the task description. In the FS scenario, the LLM was provided with three examples of posts from other users with corresponding severity ratings before assessing the new post. By comparing the results of these two approaches, we aimed to determine how much additional context in the form of examples improved the model’s ability to accurately gauge disorder severity.
FS prompting was only employed for severity evaluation because it is a more complex and challenging task compared to binary disorder detection, thereby providing a better demonstration of the potential improvement gained through FS learning.
2.1.3 Task 3: ZS Psychiatric Knowledge Assessment
This task was designed to test the LLM’s knowledge of basic psychiatric concepts. The LLM was presented with multiple-choice questions related to psychiatry. This task evaluates the LLM’s foundational understanding of psychiatric concepts and its ability to provide accurate factual information typically found in textbooks.
2.2 Datasets
We deliberately avoided datasets with automated labeling methods, such as those collecting posts from specific mental health subreddits (e.g., r/Depression) and matching them to control posts from unrelated subreddits. Additionally, we steered clear of semi/weak-labeled datasets that include posts from users who self-reported a specific disorder using phrases like "I am depressed." These approaches tend to introduce some level of unreliability, as they lack human verification to determine if a specific post or user is exhibiting mental health-related issues.
Instead, we prioritized datasets labeled by experts or crowd-sourced workers/volunteers who were trained by experts. This ensures the quality and reliability of the data used for our experiments, providing a more accurate foundation for assessing the LLM’s performance in detecting and evaluating mental health disorders.
We considered various mental disorders but ultimately focused on depression, suicide risk, and stress due to their prevalence and the relative availability of high-quality datasets. Other disorders were not included in this study due to challenges in data collection and annotation. The sensitivity of mental health data often limits the public availability of relevant social media posts, and stricter data usage agreements imposed by social media platforms [^1] The landscape of social media data collection has changed significantly in recent years. In 2023, platforms like Reddit [22] and X (formerly Twitter) [23, 24] introduced substantial costs for accessing their API s, which were previously used to create many research datasets. Furthermore, stricter data usage agreements have been implemented, making some previously popular datasets, such as the popular SMHD [25], unsharable due to updated terms of service. have further restricted access to previously used datasets. Additionally, less prevalent disorders typically have fewer human-annotated datasets available, making it difficult to compile a sufficiently large and diverse dataset for robust model evaluation.
We summarize our dataset findings in Table LABEL:tab:datasets. While not exhaustive, this table includes datasets that met our criteria for quality and relevance. However, some datasets were not directly used in this study either due to time and resource constraints or due to inaccessibility.
Table 1: Summary of Human Annotated Mental Health Datasets
Name Task Disorder(s) Source Annotator Experience Sample Size Dataset Access
Dreaddit [26] Binary Stress Classification Stress Reddit Crowdsourced 1000 posts Publicly Available
DepSeverity [27] Depression Severity Classification Depression (4 levels) Reddit Professionals 3,553 posts Publicly Available
Stress-Annotated- Dataset (SAD) [28] Stress Severity Classification Daily Stressors (9 categories) SMS-like conversations, crowdsourced data, LiveJournal Crowdsourced 6850 sentences Publicly Available
DEPTWEET [29] Depression Severity Classification Depression (4 levels) X (formerly Twitter) Crowdworkers (trained) 40,191 tweets Available Upon Request
CSSRS-Suicide [30] Suicide Risk Severity Assessment Suicide Risk (5 levels) Reddit (r/SuicideWatch) Professional psychiatrists 500 Reddit users Publicly Available
SDCNL [31] Suicide vs. Depression Classification Depression, Suicidal Ideation Reddit (r/SuicideWatch, r/Depression) N/A 1,895 posts Publicly Available
PsyQA [32] Generating Long Counseling Text Various Mental Health Disorders Mental Health Service Platform (Yixinli) Trained volunteers or professional counselors 22K questions, 56K answers Available Upon Request
MedMCQA (Psychiatry Subset) [33] Medical Question Answering Various psychiatric disorders Medical entrance exams (AIIMS & NEET PG) Questions created by exam boards 4442 (~5%) Publicly Available
PsySym [34] Symptom Identification 7 mental disorders (Depression, Anxiety, ADHD, Disorder, OCD , Eating Disorder) Reddit Trained volunteers and psychiatrists 8,554 sentences Available Upon Request
MedQA (Psychiatry Subset) [35] Medical Question Answering Various clinical topics Medical entrance exams (AIIMS & NEET PG) Exam boards 47,457 Total, ~800 psychiatry related Publicly Available
LGBTeen [36] Emotional Support and Information Queer-related topics Reddit (r/LGBTeens) Human Redditors and LLM s 1,000 posts, 11,320 LLM responses Publicly Available  [^2] While the author states that this dataset is available on Github, the dataset is not present in the Github repository provided by the author, and the author did not respond to our queries regarding the availability of the dataset
ESConv [37] Emotional Support Conversation Various Mental Disorders Crowdsourced Crowdworkers (trained) 1,053 conversations Publicly Available
Ji et al.[38] Suicidal Ideation and Mental Disorder Detection Suicidal Ideation, Depression, Anxiety, Bipolar Reddit (SuicideWatch and Mental Health-related subreddits) Experts 4800 posts Available Upon Request
Cascalheira et al. [39] Gender Dysphoria Identification Gender Dysphoria Reddit (r/GenderDysphoria) Clinicians and Students with transgender and non-binary experience 1,573 posts Available Upon Request
Hua et al. [40] Transgender and Gender Diverse Patient Identification Transgender and Gender Diverse Electronic Health Records (EHR) s Clinicians 3,350 patients Available upon request pending institutional reviews and approvals
Mukherjee et al. [41] Autism Spectrum Disorder Symptom Identification Autism Spectrum Disorder Social networks (X), special children’s organizations Researchers Not Specified Not mentioned
DATD [42] Depression and Anxiety Identification Depression and Anxiety X Not Specified 1,050 tweets Publicly Available
PAN12 [43] Sexual Predator Identification Online Grooming Perverted Justice Website, Omegle, IRC logs One Expert Hundreds of thousands of conversations Publicly Available
PJZC [44] Grooming Identification Online Grooming Perverted Justice Website, #ZIG IRC Channel, Chit-Chat Dataset Researchers 21,070 chats Publicly Available
DAIC-WOZ [45] Distress Identification Depression, Post-traumatic stress disorder (PTSD), Anxiety Interviews Experts 421 interviews Available Upon Request
D4 [46] Response Generation, Topic Prediction, Dialogue Summary, Severity Classification Depression, Suicide Risk Simulated conversations based on real patient portraits Psychiatrists, Crowdworkers 1,339 conversations Available Upon Request
DialogueSafety [47] Dialogue Safety Classification Various (Nonsense, Humanoid Mimicry, Linguistic Neglect, etc.) Chinese Online Counseling Platform, Yixinli QA Column Experienced Supporters, Counseling Psychologists, Linguistic Experts 7,935 dialogues Available Upon Request
CLPsych15 [48] Depression and PTSD Identification Depression, PTSD X Experts 1,746 tweets Available Upon Request pending institutional reviews and approvals
RED SAM [49, 50] Depression Level Identification Depression (3 levels) Reddit Crowdworkers 16,632 posts Publicly Available
While our focus was on disorder classification and evaluation, we acknowledge the existence of other relevant datasets that could evaluate LLM s’ performance on tasks like empathy, dialogue safety, emotional support provision, and counselling.
The datasets chosen for this study include:
2.2.1 Dreaddit [26]
Dreaddit is a dataset for stress analysis collected from Reddit using the PRAW API. It contains 187,444 posts from 10 subreddits across five domains: interpersonal conflict (abuse, social), mental illness (anxiety, PTSD), and financial need (financial). For annotation, posts were divided into five-sentence segments, and crowdworkers on Amazon Mechanical Turk labeled them as "Stress," "Not Stress," or "Can’t Tell." Segments with low agreement or labeled as "Can’t Tell" were discarded, resulting in a final dataset of 3,553 labeled segments. The labels indicate whether the author of the segment is expressing stress and a negative attitude towards it.
2.2.2 DepSeverity [27]
DepSeverity is a dataset designed for the identification of depression severity levels in social media posts. It contains 3,553 Reddit posts labeled across four levels of depression severity: minimal, mild, moderate, and severe. The dataset was created by relabeling an existing binary-class dataset using the Depressive Disorder Annotation scheme, which is based on six clinical resources and aligns with common depression levels in the psychological domain.
The annotation process involved two professional annotators who independently labeled each post. In cases of disagreement, a third annotator was consulted, and the majority vote was used to assign the final label.
2.2.3 SDCNL [31]
SDCNL is a dataset designed for the classification of suicidal ideation versus depression in social media posts. It contains 1,895 posts from two Reddit subreddits: r/SuicideWatch and r/Depression. The dataset was created to address the lack of research on distinguishing between these two mental health conditions, which is crucial for providing appropriate interventions.
This is the only dataset in our study that was not labeled by humans. Instead, it employed an unsupervised approach to correct initial labels based on the subreddit of origin.
2.2.4 SAD [28]
SAD (Stress Annotated Dataset) is a dataset designed for the identification of daily stressors in SMS-like conversational systems. It contains 6,850 sentences labeled across nine categories of stressors, derived from stress management literature, live conversations from a prototype chatbot system, crowdsourced data, and targeted web scraping from LiveJournal. The dataset was created to improve chatbot interactions by enabling the identification and classification of stressors in text, facilitating appropriate interventions and empathetic responses. The stress categories include: Work (1043), School (1043), Financial Problem (152), Emotional Turmoil (238), Social Relationships (99), Family Issues (101), Health, Fatigue, or Physical Pain (103), Everyday Decision Making (101), and Other (123). However, in this analysis we focused on whether stress severity without considering the source of it.
The annotation process involved crowd workers on Amazon Mechanical Turk, who labeled sentences with stressor categories and severity ratings on a 10-point Likert scale, ranging from "not stressful" (1) to "extremely stressful" (10). The dataset includes metadata such as stressor ID, sentence text, data source, labels, severity ratings, and binary fields indicating stress and COVID-19 relation. The dataset is publicly available on GitHub for research and application development purposes.
2.2.5 DEPTWEET [29]
DEPTWEET is a dataset designed for the detection of depression severity levels in social media texts. It contains 40,191 tweets labeled across four levels of depression severity: non-depressed, mild, moderate, and severe. The dataset was created by collecting tweets based on depression-related keywords and then having trained crowdworkers annotate them according to a typology based on the DSM-5 and Patient Health Questionnaire (PHQ)-9 clinical assessment tools.
The annotation process involved 111 crowdworkers who were pre-screened and trained on the labeling typology. Each tweet was annotated by at least three annotators, and the final label was determined by majority voting. A confidence score was also assigned to each label, reflecting the level of agreement among annotators. The dataset includes additional metadata such as tweet ID, reply count, retweet count, and like count.
2.2.6 MedMCQA [33]
MedMCQA is a large-scale dataset designed for medical domain question answering, focusing on a wide range of medical subjects and topics relevant to medical entrance exams. It contains 194k high-quality multiple-choice questions (Multiple Choice Question (MCQ) s) covering 2.4k healthcare topics and 21 medical subjects, including a small portion (4442) related to psychiatry. Each sample in the dataset includes a question, correct answer(s), other options, and a detailed explanation of the solution. The dataset is designed to be challenging, requiring a deep understanding of both language and medical knowledge to answer the questions correctly. It is intended to serve as a benchmark for evaluating and improving the performance of AI models in medical question answering.
2.2.7 RED SAM [49, 50]
RED SAM (Reddit Social Media Depression Levels) is a dataset designed for the detection of depression levels in social media posts. It comprises 16,632 Reddit posts labeled across three categories: "Not Depressed," "Moderately Depressed," and "Severely Depressed." The dataset was created to address the need for identifying not only the presence of depression but also its severity, which is crucial for effective treatment and intervention.
The dataset creation process involved collecting posts from relevant subreddits on mental health topics, such as "r/MentalHealth," "r/depression," "r/loneliness," "r/stress," and "r/anxiety." The collected posts were then annotated by two domain experts who followed specific guidelines to determine the level of depression reflected in each post. The annotation guidelines considered factors like the length of the post, the expression of emotions, and the presence of specific themes or language patterns.
2.2.8 CSSRS-Suicide (Excluded) [30]
CSSRS is a dataset designed for assessing the severity of suicide risk in individuals based on their social media posts. It contains 500 Reddit users from the r/SuicideWatch subreddit, labeled across five levels of suicide risk severity: Supportive, Indicator, Ideation, Behavior, and Attempt. The dataset was created to enable early intervention and prevention of suicide by identifying individuals who may be at risk.
The annotation process involved four practicing clinical psychiatrists who followed the guidelines outlined in the Columbia Suicide Severity Rating Scale (C-SSRS). The dataset includes the Reddit posts of the 500 users, along with their corresponding suicide risk severity labels.
We originally planned to include the CSSRS-Suicide dataset in our analysis, However, as the dataset is made up of posts from 500 people, with the number of posts included for each user varying greatly, from 1 in the lower case and more than 100 in the upper case. This made it difficult to evaluate the dataset fairly, with some users including more information than others. Furthermore, evaluating the users with too many posts would prove quite costly and exceed the context window of many of the smaller LLM s tested. Upon trying to include only one post per user (e.g. by selecting only the first post for each user), the results were significantly worse and as such the dataset was discarded.
2.3 Sampling
As all testing was conducted through API s, and for the bigger models, such as GPT-4, the costs could escalate rapidly, we employed two strategies for cost saving.First, we utilized the test dataset only if it was pre-split. Second, if the dataset exceeded 1000 examples, we employed a fair random sampling of 1000 instances. This sample size was chosen based on Hoeffding’s inequality (Equation 1) [51], which provides a general upper bound for the error difference between a sample and the entire population, given a random sample of that size.
Datasets containing severity metrics like DepSeverity and SAD were binarized by assigning a "False" label (0) to posts/users with the minimum severity score and a "True" label (1) to those with any higher severity score.
$P\left(\left|\nu-\mu\right|>\varepsilon\right)\leq 2e^{-2\varepsilon^{2}N}$ (1)
* $\nu$ is the empirical mean (sample average) * $\mu$ is the true mean (population average) * $\varepsilon$ is the deviation from the mean * $N$ is the number of samples
Plugging in the values $N$ = 1000 and $\varepsilon$ = 0.05 into Equation 1 (representing a maximum discrepancy of 5%), we calculate that the probability of the in-sample and out-of-sample error deviating by more than 5% is merely 0.0135. This implies a greater than 98.5% probability that the error will remain within a 5% range of the true error.
2.4 Models
To ensure a comprehensive assessment of LLM capabilities across diverse architectures and sizes, we selected a wide range of models for our experiments. We included both large, state-of-the-art models known for their performance on various natural language processing tasks, as well as smaller, more accessible models that could potentially be deployed in resource-constrained environments.
Our model selection encompassed both closed-source models, accessible through APIs, and open-source models, which offer greater transparency and flexibility for research purposes. We included models from various families, including OpenAI’s GPT series, Anthropic’s Claude models, Mistral AI’s models, Google’s models, Meta’s Llama series, and finally, Phi-2 and MentaLlama, a model specifically fine-tuned on mental health-related data. Additionally, we attempted to include Gemini 1.5 Pro in our evaluation however due to rate limit constraints, and challenges mentioned in 5, we were not able to include it.
Table 2 provides a summary of the models used in our experiments, along with their respective sizes and the API s through which we accessed them.
Table 2: Summary of large language models evaluated in this analysis
Model Size (Parameters) Source API Service Provider
Claude 3 Haiku [52] Proprietary Anthropic Claude API
Claude 3 Sonnet [52] Proprietary Anthropic Claude API
Claude 3 Opus [52] Proprietary Anthropic Claude API
Claude 3.5 Sonnet [53] Proprietary Anthropic Claude API
Gemma 2b [54] 2B Google Nvidia NIM
Gemma 7b [54] 7B Google Groq
Gemma 2 9b [55] 9B Google Nvidia NIM
Gemma 2 27b [55] 27B Google Nvidia NIM
Gemini 1.0 Pro [8] Proprietary Google Gemini API
GPT-4o mini [56] Proprietary OpenAI OpenAI API
GPT-3.5-Turbo [57] 175B OpenAI OpenAI API
GPT-4o [58] Proprietary OpenAI OpenAI API
GPT-4-Turbo [7] Proprietary OpenAI OpenAI API
GPT-4 [7] Proprietary OpenAI OpenAI API
Llama 3 8b [59] 8B Meta Groq
Llama 3.1 8b [59] 8B Meta Together AI
Llama 2 13b [60] 13B Meta Together AI
Llama 2 70b [60] 70B Meta Together AI
Llama 3 70b [59] 70B Meta Groq
Llama 3.1 70b [59] 70B Meta Together AI
Llama 3.1 405b [59] 70B Meta Together AI
Mistral 7b [61] 7B Mistral AI Together AI
Mistral NeMo [62] 12B Mistral AI Mistral AI API
Mixtral 8x7b [63] 45B (13B) Mistral AI Groq
Mixtral 8x22b [63] 141B (39B) Mistral AI Together AI
Mistral Medium [64] Proprietary Mistral AI Mistral AI API
Mistral Large [65] Proprietary Mistral AI Mistral AI API
Mistral Large 2 [66] Proprietary Mistral AI Mistral AI API
Phi-2 [67] 2.7b Phi-1 Labs Nvidia NIM
Phi-3-mini [68] 3.8B Phi-1 Labs Nvidia NIM
Phi-3-small [68] 7B Phi-1 Labs Nvidia NIM
Phi-3-medium [68] 14B Phi-1 Labs Nvidia NIM
Qwen 2 72b [69] 72B Alibaba Cloud Together AI
Notes: The Mixtral 8x7b and 8x22b models have two sizes because the total size and the number of parameters utilized at any point in inference are different.
For some of the side experiments discussed in Section 4, a subset of the available models was utilized due to cost constraints. However, the selected subset was carefully chosen to ensure representation from all relevant model families, maintaining the diversity of the study.
2.5 Prompt Templates
The prompt, or instruction, given to an LLM plays a great role in its performance on a given task. Prior research has shown that even slight variations in prompt wording can significantly impact the quality of LLM outputs [70]. Therefore, we dedicated considerable effort to crafting effective prompts tailored to each task in our study.
To optimize LLM performance on social media post analysis, we employed prompts that encouraged the models to adopt the role of a psychiatrist. Role-playing has proven effective in eliciting relevant and contextually appropriate responses from LLM s [71, 72]. We also utilized the concept of prompt templates, which consist of a fixed structure with variable elements that can be adapted for different tasks. This approach ensured consistency across experiments while minimizing prompt-based bias.
We developed multiple prompt templates for each task category: binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment. These templates varied in their structure and level of explicit instruction, aiming to accommodate the diverse capabilities of the LLM s included in our study.
2.5.1 Binary Disorder Detection Prompt Templates
In the ZS binary disorder detection task, we employed four distinct prompt templates, designated BIN-1 to BIN-4. By utilizing multiple prompt templates, we aimed to assess the impact of both open-ended and structured prompts on the performance of various tested LLM s. This approach allowed us to evaluate LLM s’ sensitivity, variability, and adaptability to diverse prompting styles, highlighting their strengths and weaknesses in different scenarios.
As shown in 3, Prompts BIN-1 and BIN-2 utilized a straightforward structure, first instructing the LLM to assume the role of a psychiatrist and then directly asking whether the poster of a given social media post exhibited a specific disorder. We included a final instruction to elicit a "yes" or "no" response for easier parsing of the LLM output.
Table 3: Task 1 (Binary Disorder Classification) Prompt Templates
Prompts BIN-1 and BIN-2
As a psychologist, analyze the provided social media post(s).
Is the poster depressed?/ Does the poster exhibit clear symptoms of depression?
(Answer with Yes or No only without explaining your reasoning)
Prompt BIN-3
[Task] Act as a psychologist. Label the writer of the provided social media post as ’depressed’ or ’not depressed’ according to provided guidelines. [Guidelines] - Concise Response: Respond only with ’Yes’ (depressed) or ’No’ (not depressed). - No Explanations: Don’t provide explanations for your assessment. - Ambiguity: If the post is unclear, choose the most probable label. [Post]
Prompt BIN-4
Analyze the following social media post to determine if the writer exhibits clear
symptoms of depression according to provided guidelines.
… Same as BIN-3
Prompts BIN-3 and BIN-4 introduced a more structured format, explicitly separating the prompt into "Task," "Guidelines," and "Post" sections. The guidelines emphasized the desired output format (yes/no), discouraged explanations, and instructed the LLM to choose the most probable label in cases of ambiguity. This structure aimed to improve the LLM s’ adherence to the task instructions and reduce the incidence of irrelevant or unparseable responses. The inspiration for this structure was from Gemini, as this is the structure used for its training [8].
2.5.2 Disorder Severity Evaluation Prompt Templates
For the disorder severity evaluation task, we iteratively refined the prompt template based on initial observations from the binary task and the performance of various LLM s. We began with a baseline prompt, SEV-1, which was a modified version of the binary prompt template BIN-1. The key modification in SEV-1 was the addition of a severity scale, representing varying levels of disorder severity. Figure 1 illustrates all four Severity prompts (SEV-1 to SEV-4).
During the design of this initial prompt, we observed that most LLM s performed better when provided with the full context of each severity level, rather than receiving no context or only the extreme values (0 and 4).
Subsequent refinements aimed to address the tendency of some expensive models to provide verbose explanations despite being instructed to give concise answers. We systematically introduced modifications to the prompt, resulting in three additional versions: SEV-2, SEV-3, and SEV-4.
  1. 1.
    SEV-2: Modified the presentation of the severity scale from free text to a numbered list to enhance clarity.
  2. 2.
    SEV-3: Incorporated an ambiguity clause (similar to the one used in BIN-3 and BIN-4) to guide responses when the post was unclear.
  3. 3.
    SEV-4 :Reiterated the instruction to provide a concise answer without explanation. This simple repetition proved surprisingly effective, especially for Mistral models. The added line is henceforth referred to as repetition line
Figure 1: Severity Task Prompts
To evaluate the effectiveness of these modifications, we conducted experiments on a sample of 1000 posts from the DEPTWEET dataset, comparing the performance of different LLM s across the various prompt versions (P1-P4). The results, presented in Table 4, demonstrate a general trend of improved performance from P1 to P4, particularly for models like Gemini 1.0 Pro and Llama 3 70b. Notably, even models that did not show significant improvement experienced only minor decreases in performance. Hence, prompt SEV-4 was chosen as the main prompt for the severity task.
Table 4: Performance metrics and invalid responses of different models across the various severity prompts
Model/Prompt P1 P2 P3 P4
GPT-3.5-Turbo MAE: 0.713 MAE: -0.001 MAE: -0.029 MAE: -0.066
BA: 43.6% BA: -1.4% BA: -1.7% BA: -0.4%
IR: 0 IR: 55 IR: 0 IR: 0
Gemini 1.0 Pro MAE: 0.629 MAE: -0.039 MAE: -0.095 MAE: -0.133
BA: 42.7% BA: +2.8% BA: +7.6% BA: +11.5%
IR: 0 IR: 0 IR: 0 IR: 0
Claude 3 Haiku MAE: 0.97 MAE: -0.041 MAE: -0.069 MAE: -0.092
BA: 29.8% BA: +2.3% BA: +3.2% BA: +4.5%
IR: 70 IR: 22 IR: 18 IR: 4
Mistral Medium MAE: 0.608 MAE: +0.001 MAE: -0.051 MAE: -0.052
BA: 46.2% BA: +2.2% BA: +5.8% BA: +5.7%
IR: 38 IR: 39 IR: 3 IR: 0
Gemma 7b MAE: 0.892 MAE: +0.027 MAE: +0.058 MAE: +0.031
BA: 35.5% BA: -1% BA: -1.9% BA: -1.3%
IR: 4 IR: 4 IR: 7 IR: 0
Llama 3 8b MAE: 1.093 MAE: +0.146 MAE: -0.043 MAE: -0.324
BA: 32% BA: -5.7% BA: -1.3% BA: +5%
IR: 137 IR: 117 IR: 57 IR: 7
Llama 3 70b MAE: 0.949 MAE: -0.062 MAE: -0.105 MAE: -0.275
BA: 29% BA: +4.3% BA: +6.0% BA: +13.3%
IR: 23 IR: 22 IR: 17 IR: 0
Notes: BA: Balanced Accuracy, MAE: Mean Absolute Error, IR: Invalid Responses
For the FS disorder severity evaluation task, the SEV-4 prompt was initially used as a foundation. Subsequently, three hand-selected examples were Added to the prompt as a guidance (The chosen examples are not sent to the LLM s for inference). Table 6 clearly details these prompt modifications.
Table 5: Prompt SEV-4 (FS)
Prompt
SEV-4 (ZS) prompt contents
Some examples for guidance:
Example 1:
Post: <post contents>
Response: Yes/No
Example 2:
Post: <post contents>
Response: Yes/No
Example 3:
Post: <post contents>
Response: Yes/No
The post to be analyzed is:
Table 6: Prompt KNOW-1
Prompt KNOW-1
As a psychologist, please analyze the following MCQ question and select the most appropriate answer.
Do not provide any explanation for your choice.
Indicate your choice with a single English letter.
(Provide only one answer)
2.5.3 Psychiatric Knowledge Assessment Prompt Template
For the psychiatric knowledge assessment task, we employed a simple prompt template (KNOW-1) that instructed the LLM to answer a multiple-choice question with only the corresponding letter (A, B, C, or D) and no explanation. There was no need for any iterative experimentation. The prompt is illustrated in Table 6
2.6 Evaluation Metrics
To assess the performance of the LLM s across our defined tasks, we primarily employed accuracy as our evaluation metric. For most tasks, we applied fair random sampling to the datasets, ensuring balanced class distributions. This balanced sampling naturally leads to the calculation of Balanced Accuracy (BA), which is the average of recall obtained on each class. BA is particularly well-suited for classification tasks where class imbalance may be present, as it provides a more equitable assessment of the model’s ability to correctly identify both positive and negative instances compared to traditional accuracy.
However, in cases where fair random sampling was not feasible due to limitations in the dataset, such as the SAD dataset with its sparse distribution of severity classes, we utilized regular random sampling with a random state of 42 and evaluated performance using traditional accuracy. Similarly, for the MedMCQA dataset, which involves multiple-choice question answering without distinct class labels, we also relied on accuracy as the primary evaluation metric.
For the disorder severity evaluation task, we employed two metrics to assess the performance of LLM s: BA and Mean Absolute Error (MAE). BA, in this context, refers to the proportion of posts for which the LLM predicted the exact severity level assigned by human annotators. It serves as a measure of how often the model gets the severity classification exactly right. On the other hand, MAE quantifies the average magnitude of the errors in the predicted severity ratings. It provides a complementary perspective on model performance by assessing how close, on average, the predicted severity levels are to the true levels, even if they are not exactly correct. A model could potentially improve in one metric while worsening in the other.
2.7 Parsing Model Outputs
Parsing, the process of interpreting and extracting meaningful information from LLM outputs, proved to be a non-trivial aspect of our methodology. While seemingly straightforward for tasks requiring simple responses like "yes/no" or numerical ratings, the reality was more complex. LLM s, particularly larger and more sophisticated models, often deviate from explicit instructions regarding output format, frequently providing explanations or incorporating the requested answer within a longer response.
We observed that model adherence to instructions varied widely and did not necessarily correlate with model size. For instance, smaller models like Llama 3 8b, as well as larger ones like Llama 3 70b and GPT-4, demonstrated a high degree of compliance with instructions to output only the answer. Conversely, other models, such as the 2 billion parameter Phi-2 and the seemingly large Mistral Large and Claude 3 Opus, frequently struggled to adhere to the desired format, often opting to provide reasons and explanations for their provided label.
To address these challenges, we developed custom parsers for each task:
  1. 1.
    Binary Disorder Detection: Our parser searched for the text "yes" or "no" in the LLM output. If only one option was found, it was taken as the model’s answer. If both were present, or neither was found, the output was considered to be an invalid response.
  2. 2.
    Disorder Severity Evaluation: The parser searched for numerical values within the specified severity range. If a single valid number was found, it was accepted as the answer. If multiple numbers were found, the number was outside the valid range, or no number was found, then output was considered to be an invalid response.
  3. 3.
    Psychiatric Knowledge Assessment: Parsing multiple-choice answers proved more challenging, particularly for the letter "a," which could be either the answer or part of a word like "an." We employed an iterative approach, refining the parser’s regular expressions by examining invalid outputs. This iterative process revealed that LLM s tend to follow a limited set of patterns in their responses, making it feasible to achieve reliable parsing with relatively few rules. Interestingly, only eight rules were required to parse the majority of outputs across all LLM s.
2.8 Additional Considerations
In addition to the evaluation metrics and parsing strategies discussed previously, there are a few additional considerations worth noting.
First, despite our efforts in prompt engineering, certain LLM s, such as Mistral Medium and Claude 3 Opus, consistently struggled to adhere to the instructions provided in the prompts across various tasks. To mitigate this issue, we employed modified prompts in some experiments, primarily by incorporating the "repetition line" from prompt SEV-4, which explicitly reiterated the instruction to provide a concise answer without explanation. This modification generally improved the adherence of these models to the desired output format.
Second, our analysis included several smaller, specialized experiments designed to evaluate specific aspects of LLM performance or explore particular research questions. Due to their focused nature, the details of these experiments, including their implementation, results, and discussion, are presented separately in Section 4 to maintain clarity and coherence in the main body of our analysis.
Third, during evaluation, we compared two approaches: setting a maximum output token limit of 2 tokens versus allowing the models to respond freely without truncation. The rationale behind the former approach was that it would likely include the keyword for the diagnosis we were seeking. Conversely, for models that did not adhere well to prompt instructions, the diagnosis might appear later in their response. Our findings revealed that the longer a model’s response, the more likely it was to become invalid due to the possibility of it using ’yes’ or ’no’ when providing explanations for its decision. Consequently, we decided to impose a response length limit of 2 tokens for all models tested. Fig 2 presents a summary of each model’s invalid response percentage.
The GPT family showed strong performance, with an average invalid response rate of 0.45%. GPT-4 demonstrated impressive performance, with a reduced invalid % rate to 0.02%. However, this positive trend reversed with GPT-4-Turbo, where the rate increased to 1.64%. The subsequent GPT-4o and GPT-4o mini versions returned to low invalid rates of 0.22% and 0.0%, respectively.
As for the Llama family, Llama 3 represented a marginal improvement over Llama 2’s already impressive invalid percentage numbers (0.43% to 6.37%). However, the most recently released Llama 3.1 models perform significantly worse potentially due to their ethical guards [59]. Llama 3.1 70b is not affected as much, showing a 0.42% worse performance than Llama 3 70b, which had a 0.37% invalid rate. However, the other two models, and especially the Llama 3.1 405b model, are significantly worse, with the Llama 3.1 405b reaching a 9.0% invalid rate.
The Mistral family showed steady improvement in following instructions, starting with Mistral 7b at 12.29% and decreasing to nearly 0% for both Mistral Nemo 7b and Mistral Large 2. Despite this overall positive trend, Mistral Large unexpectedly had a high invalid response rate of 19.68%, which was later corrected in Mistral Large 2, bringing the rate back down to 0%.
The Gemma family displayed a stable pattern in invalid response rates, with most models performing consistently across iterations. The notable exception was Gemma 2b, which, likely due to its smaller size, recorded the highest invalid rate at 33.53%. Surprisingly, Gemma 7b maintained a low rate of 1.88%, similar to the larger models like Gemma 1.0 Pro (1.37%), Gemma 2 9b (0.52%), and Gemma 27b (0.35%).
The Phi, Qwen, and Claude families exhibited varied performances, with the Phi models showing a range of invalid response rates from 0.07% to 27%, with Phi-3-medium being the best. The Qwen 2 72b model had a moderate invalid response rate of 1.46%. The Claude family displayed significant variability, with rates ranging from 0.8% to 20.49%. Surprisingly Claude 3 Opus showed worse performance compared to the significantly cheaper Clause 3 Haiku, with both being significantly worse than Claude 3 Sonnet.
2022-11-282023-03-142023-07-182023-09-272023-11-142023-12-092023-12-122023-12-232024-02-212024-02-262024-03-142024-04-182024-05-132024-06-062024-06-202024-06-272024-07-022024-07-182024-07-232024-07-240.010.010.010.010.10.10.10.1111110101010100100100100Claude 3 HaikuClaude 3 SonnetClaude 3.5 SonnetClaude 3 OpusGemma 2bGemma 7bGemma 2 9bGemma 2 27bGemini 1.0 ProGPT-4o miniGPT-3.5-TurboGPT-4oGPT-4-TurboGPT-4Llama 3 8bLlama 2 13bLlama 2 70bLlama 3 70bLlama 3.1 8bLlama 3.1 70bLlama 3.1 405bMistral 7bMistral NeMoMixtral 8x7bMixtral 8x22bMistral MediumMistral LargeMistral Large 2Phi-2Phi-3-miniPhi-3-smallPhi-3-mediumQwen 2 72bRelease Dateof Invalid Responses (Log Scale)
Figure 2: Invalid Response % Statistics Sorted by model release date
3 Results & Discussion
In this section, we present the results of our experiments, analyzing the performance of various LLM s across the three tasks defined in Section 2. We begin by examining the performance variability between runs across different LLM s using consistent parameters for all experiments. This analysis establishes a foundation for assessing the reliability and reproducibility of the presented results, determining whether observed improvements between prompts are valid and significant or merely due to chance. Next, we delve into the results of the ZS binary disorder detection task (Section 3.3), followed by the ZS/FS disorder severity evaluation task (Section 3.4), and conclude with the ZS psychiatric knowledge assessment task (Section 3.5).
In our evaluation of the LLM s across the three distinct tasks, we observed a wide range of performance capabilities and several noteworthy trends. In the ZS binary disorder detection task, GPT-4, GPT-4o, Mistral NeMo and Llama 2 70b consistently outperformed other models. Additionally, the choice of prompt significantly influenced model performance, with some models favoring more structured prompts while others performed better with open-ended ones. Notable mentions in this category include the Gemma 7b model which achieved 76.70% (BIN-4) accuracy on Dreaddit, which is significantly higher than the previous BIN-1 best of 74.50% of the Mistral NeMo, Mistral NeMo for competing with models much larger than it in size.
The ZS/few-shot disorder severity evaluation task revealed that most models benefited from few-shot learning, where they were provided with a few examples before making predictions. However, the degree of improvement varied across models and datasets, highlighting the need for careful consideration of both model architecture and data characteristics. Finally, in the ZS psychiatric knowledge assessment task, we found that the more recent models generally performed better, compared to older, much larger models (e.g. Llama 3 70b and Phi-3-medium competing with GPT-4, GPT-4-Turbo, Claude 3.5 Sonnet, and others). However, with models released around the same time period, model size seemed to mostly be the deciding factor (Llama 3.1 models are excluded from this finding due to their tendency to filter/not respond to our prompts due to their sensitive mental health nature).
3.1 Performance Variability Experiment
In the performance variability experiment, we evaluate the inherent performance variability of each LLM on the same task. The experiment involved repeatedly evaluating each model on the same dataset five times with the standard parameters we typically use (e.g., temperature 0). This approach allows us to approximate the natural performance fluctuations, enabling us to judge whether any observed performance enhancement brought about by changes to the prompt or other factors is a meaningful improvement or a random fluctuation due to the model’s inherent variability. The experiment was conducted on 1000 fairly sampled instances from the DEPTWEET dataset.
3.2 Performance Variability Experiment
In the performance variability experiment, we evaluate the inherent performance variability of each LLM on the same task. The experiment involved repeatedly evaluating each model on the same dataset five times with the standard parameters we typically use (e.g., temperature 0). This approach allows us to approximate the natural performance fluctuations, enabling us to judge whether any observed performance enhancement brought about by changes to the prompt or other factors is a meaningful improvement or a random fluctuation due to the model’s inherent variability. The experiment was conducted on 1000 fairly sampled instances from the DEPTWEET dataset.
Table 7: Performance Variability on DEPTWEET Across Five Runs
Model Standard Deviation (%)
Gemma 2 9b 0.05
Gemma 2 27b 0.1
GPT-3.5-Turbo 0.4
Llama 2 70b 0.2
Llama 3 70b 0.1
Mixtral 8x22b 1.6
Phi-3-small 0.05
Phi-3-medium 0.16
Qwen 2 72b 0.04
The results in Table 7 show that most models exhibited performance fluctuations of less than 0.5%, with the exception of Mixtral 8x22b, which showed a 1.6% deviation. Models not listed in the table had a standard deviation of 0%, indicating no performance variability across the five runs. These results suggest that observed changes in model performance (as long as they are larger than the observed fluctuations) are likely due to our modifications (e.g., to the prompt utilized) rather than inherent model variability.
3.3 Binary Disorder Classification (Task 1)
In the ZS binary disorder detection task, we evaluated the performance of various LLM s in identifying the presence or absence of depression, suicide risk, and stress in social media posts. We began by assessing their performance using the BIN-1 prompt template, followed by an analysis of the impact of modifications introduced in prompts BIN-2, BIN-3, and BIN-4.
3.3.1 Results of Prompt BIN-1
An analysis of Figure 3 reveals that the performance of each model varies across datasets. However, a general trend emerges: OpenAI stands out with three of the six best-performing models. Specifically, GPT-4 achieves top scores of 85.20% on the SAD dataset and 85.00% on the DEPTWEET dataset, while GPT-4o excels with 71.20% on the SDCNL dataset. In addition to OpenAI’s success, Llama 2 70b performs best in the DepSeverity dataset with 73.58%, Mistral NeMo leads in both Dreaddit Test and RED SAM with 74.50, and 66.00% respectively.
OpenAI models consistently demonstrate strong performance. GPT-4 achieves the highest accuracy among the tested models, reaching approximately 85% on both the DEPTWEET and SAD datasets. GPT-4-Turbo and GPT-4o generally follow closely behind GPT-4, though with some performance fluctuations. GPT-4-Turbo typically lags slightly, outperforming GPT-4 in two instances (63.7% in RED SAM vs. 62.1%, and 70.8% in SDCNL vs. 69.3%). GPT-4o exhibits more significant fluctuations, sometimes exceeding all other variants (e.g., 71.2% in SDCNL) but also deviating from GPT-4 by up to 9.8% in SDCNL. GPT-3.5-Turbo consistently follows closely behind GPT-4, deviating by approximately 5% at worst, which is notable considering its much smaller size. The newest and most cost-efficient model, GPT-4o mini, generally performs worse compared to GPT-3.5-Turbo.
The Llama 2 70b model demonstrates notable performance, achieving 73.58% in DepSeverity, unexpectedly outperforming the Llama 3 70b (70.5%) and even surpassing GPT-4, which achieves 72.6%. The newer Llama 3 models exhibit greater consistency, with deviations of only 3-10% between the 8b and 70b variants, compared to 10-20% deviations in the Llama 2 family. However, the Llama 3.1 family of models was a significant disappointment, with the Llama 3.1 8b model performing considerably worse compared to the earlier version, except in two datasets (SAD & DEPTWEET), where the performance is generally comparable. A similar observation applies to the Llama 3.1 70b model, which is generally closer in performance to the earlier version but actually outperforms the Llama 3 70b model in the SAD and DEPTWEET datasets. Lastly, the Llama 3.1 405b model was a major disappointment, performing even worse than the Llama 3.1 70b model across all six datasets.
Mistral 7b shows moderate performance, achieving an accuracy of 55% (± 2.8%) across datasets, barely surpassing the baseline accuracy of a random classifier. Although it outperforms Llama 2 13b, it is surpassed by Llama 3 8b. The Mistral Medium model demonstrates similar performance to Mistral 7b. Mistral Large initially encounters difficulties due to instruction-following issues but shows significant improvement with the BIN-1.1 prompt, achieving a 5-25% increase in accuracy depending on the dataset, with the exception of Dreaddit. The Mistral NeMo model stands out as the best-performing model on the Dreaddit Test dataset with 74.5%, outperforming GPT-4’s 72%. Beyond Dreaddit, Mistral NeMo performs better than Mistral 7b, but is generally outpaced by Llama 3 8b. As for Mistral Large 2, its performance is much better than Mistral Large 1, even mostly outperforming Mistral Large 1 with the BIN-1.1 prompt. However, it performs worse overall compared to the much smaller Mistral NeMo model
The Claude family presents varied results. Claude 3 Haiku often performs similarly to or even surpasses Claude 3 Sonnet, particularly in Dreaddit (56.6% vs. 54.2%) and SAD (74.6% vs. 71.3%). Claude 3 Sonnet shows a significant advantage over the Haiku model only in DEPTWEET (81.5% vs. 75.5%). Claude 3 Opus, Anthropic’s flagship model, also struggles initially (due to the same instruction-following problems) but improves with BIN-1.1, though only marginally outperforming Haiku and Sonnet in most cases, except SAD (81.3% vs. 74.6% and 71.3%). The newer Claude 3.5 Sonnet model outperforms all of its siblings, including the Claude 3 Opus model (even when using the BIN-1.1 prompt) in all but two instances (78.5% vs 81.3% in SAD, and 58.1% vs 58.4% in RED SAM). Overall, the Claude family demonstrates inconsistency in model performance compared to other families. However, the Claude 3.5 Sonnet model is overall the best-performing model from this family.
Gemini 1.0 Pro generally trails GPT-3.5-Turbo, only slightly outperforming it in the DEPTWEET dataset. This performance is notable given GPT-3.5-Turbo’s proximity to GPT-4’s capabilities. Gemma 7b performs similarly to Mistral 7b, while Gemma 2b and Phi-2 lag significantly, even falling below baseline performance, with Gemma 2b reaching only 5% and 13% accuracy in SDCNL and DepSeverity, respectively.
Looking at the newest model families—Gemma 2 and Phi-3—there is clear inconsistency in the results obtained between the Gemma 2 9b and the Gemma 2 27b models. In 5 out of 6 instances, the Gemma 2 9b model outperforms its larger sibling, with the largest difference being in the DEPTWEET dataset (80.4% vs 77.9%). Overall, the two models are close in performance; however, it is unexpected that the smaller model is consistently more performant.
As for the Phi-3 family, we observe a similar trend of inconsistency. The Phi-3-mini model consistently outperforms the larger Phi-3-small model in 5 out of the 6 datasets, and even outperforms the Phi-3-medium model in the RED SAM dataset (57.8% vs 57%). The Phi-3-medium model generally outperforms its smaller siblings except for two instances.
Lastly, the Qwen 2 72b model performs well across most datasets without any specifically noteworthy performance.
Figure 3: Results of Prompt BIN-1 on Task 1
Figure 4 summarizes model performance deviations, indicating how far each model is from the best-performing model for each dataset. This allows for an easy comparison between models and provides a general idea of their overall performance. A clear observation is that the graph is separated into clusters of similarly performing models.
To start, GPT-4 excels with an exceptionally low average deviation of 1.55%. This places it at the top of the performance spectrum. Following closely behind, we have a cluster of high performers, including GPT-4-Turbo, GPT-3.5-Turbo, Llama 3 70b, Llama 3.1 70b, Llama 2 70b, GPT-4o, Claude 3.5 Sonnet, Gemini 1.0 Pro, Qwen 2 72b, and Mistral NeMo, with deviations ranging from 4.40% to 6.60%. It is noteworthy that the GPT-3.5-Turbo model slightly outperforms the GPT-4o model (4.58% vs. 5.66% average deviation), while being only very slightly outperformed by the GPT-4-Turbo model (4.40%). This highlights the exceptional performance of the older model.
Next, models such as Mistral Large 2, Claude 3 Opus, and Mistral Large, both using the BIN-1.1 prompt, along with Phi-3-medium, demonstrate similar performance with an average deviation of around 8% (± 0.3%). Considering the Phi-3-medium model’s relatively small size of 14 billion parameters, its performance is particularly impressive and indicates a high level of optimization.
The next cluster includes moderately performing models with average deviations ranging from 9.28% to 11.73%. This cluster includes GPT-4o mini, Gemma 2 9b/27b, Claude 3 Haiku/Sonnet, Phi-3-mini, Mixtral 8x7b, and Llama 3 8b.
The remaining clusters include models with average deviations of more than 12%, such as Mistral Medium, Phi-3-Small, Mistral Large, and Claude 3 Opus (with the prompt BIN-1), Gemma 7b, Mistral 7b, Llama 2 13b, Phi-2, Gemma 2b, and others. Most disappointing in this category are the Llama 3.1 405b and, to a much lesser extent, the Llama 3.1 8b, which performs poorly despite being newer and of much larger models.
Overall, several models stand out for their notable (whether positive or negative) performance on the six binary datasets. OpenAI stands out for its performance once more, with the top three models belonging to it, and the next three belonging to Meta. Sadly, Llama 3.1 70b performs slightly worse than the older Llama 3 70b model, but this is much better than the Llama 3.1 405b model, which was a major disappointment.
Mistral NeMo slightly outperforms the Mistral Large 2 model despite their significant size difference (14b vs. 123b). Phi-3-medium and Gemma 2 9b also perform admirably, holding their own against much larger models. Overall, Mistral NeMo emerges as the top model in the sub-14 billion parameter category, followed by Phi-3-medium. The Llama 3 70b model leads in the sub-70 billion parameter category, followed closely by the Llama 3.1 70b model. The Phi-3-mini model also stands out, performing comparably to Claude 3 Haiku/Sonnet despite being only 3.8 billion parameters in size.
Figure 4: Deviation of each model’s accuracy from the best model’s accuracy on Task 1 using Prompt BIN-1
3.3.2 Results of Prompts BIN-2/3/4
Table 8 presents the changes in model performance when moving from prompt BIN-2 to BIN-3, and BIN-4 respectively. Comparing the improvement when going from prompt BIN-1 to prompt BIN-2, we can observe the effect of changing the main task from checking whether the poster suffers from a disorder (i.e., ’is the user stressed’) to checking whether the poster shows or exhibits symptoms of a disorder (’Does this poster exhibit clear symptoms of stress’). This rephrasing results in different outcomes, specifically creating a less strict boundary for a true (i.e., stressed) label.
Models such as Gemini 1.0 Pro, Mistral Medium, Mistral Large 2, Llama 2 13b, Llama 3.1 8b/70b/405b, Mixtral 8x7b/22b, Mistral 7b, Gemma 7b, Gemma 2 9b/27b, Phi-3-mini/small, and Qwen 2 72b exhibited improvements moving from prompt BIN-1 to BIN-2, particularly on the Dreaddit and DepSeverity datasets. These improvements were more pronounced in models with initially lower performance, generally stabilizing around the 65-70% accuracy range. For instance, Gemini 1.0 Pro, which started with an accuracy above 60%, improved to 72.20% on the Dreaddit Test. Mixtral 8x22b, Gemma 7b, and Mistral Large 2 demonstrated substantial improvements, nearing the 70% accuracy mark, with increases of 12.29%, 13.68%, and 14.80%, respectively. This performance approaches that of the previously best-performing model, Mistral NeMo, which had achieved 74.50% accuracy on the Dreaddit Test with BIN-1. The Phi-3-mini and Phi-3-small models showed moderate improvements of 4.80% and 2.30%, respectively, while Qwen 2 72b improved by 5.70%. In contrast, the SDCNL dataset generally did not exhibit similar improvements, potentially due to lower labeling quality, which may have hindered the models’ ability to generalize improvements.
When examining the transitions from BIN-1 to BIN-3 and BIN-4, fewer models exhibited improved performance. Notably, models such as GPT-3.5-Turbo, Claude 3 Haiku, Llama 3 8b, Llama 2 70b, Llama 3.1 405b, Mistral NeMo, Phi-3-medium, and Qwen 2 72b showed declines ranging from 2% to over 20%, depending on the model and dataset. This trend indicates a general preference for the BIN-2 prompt, likely due to its clearer less structured approach.
Among the models showing notable improvements, Gemini 1.0 Pro demonstrated significant gains on Dreaddit and DepSeverity, with increases of +8% and +4% (BIN-4), respectively. Gemma 7b also excelled, achieving a 22.58% increase on Dreaddit with prompt BIN-4, reaching 76.70% accuracy—surpassing the previous best of 74.50% achieved by Mistral NeMo. On the DepSeverity dataset, Gemma 7b also improved significantly with a gain of 11.88% (BIN-3). Gemma 2b, while not showing a clear preference between BIN-3 and BIN-4, exhibited fluctuations with significant improvements, notably a +45% increase (BIN-4) on SDCNL, starting from an anomalous 5%, and a +51% increase (BIN-3) on DepSeverity from a baseline of 13%.
Llama 2 13b displayed a strong preference for BIN-4 over BIN-3, showing improvements with BIN-4 of 9.41% on Dreaddit and 13.80% on DepSeverity, as opposed to the significant performance degradation with BIN-3 (-42%, -26%). Meanwhile, the Llama 3.1 8b model showed substantial improvements with both BIN-3 and BIN-4, particularly with increases of +3.7% (BIN-3) and 10.80% (BIN-4) on Dreaddit, and a +6.8% gain (BIN-4) on DepSeverity. The Llama 3.1 70b model showed minor improvements of approximately 1-2% with BIN-4 across datasets.
The Mistral and Mixtral models consistently improved with each prompt, with a clear preference for BIN-4. Mistral Medium achieved a notable 15.40% improvement on DepSeverity with BIN-4, reaching a total accuracy of 75.10%. Mixtral 8x7b and Mixtral 8x22b also showed substantial improvements, with increases of 10.75% on Dreaddit and 21% on DepSeverity, respectively. Mistral Large 2 similarly preferred BIN-4, improving by 10.40% on Dreaddit and 2.50% on DepSeverity. In contrast, Mistral NeMo exhibited declines with both BIN-3 and BIN-4, showing significant decreases in performance on Dreaddit and DepSeverity.
The Phi-3 models displayed mixed results, with Phi-3-mini showing a slight improvement of 0.80% on Dreaddit with BIN-4, but declines in other metrics. Phi-3-small exhibited moderate improvements with BIN-2 and BIN-3, and a notable 5.50% increase with BIN-4. Phi-3-medium, however, showed declines across all prompts.
Gemma 2 9b and Gemma 2 27b displayed consistent improvements, particularly with BIN-4 across all datasets. The highest accuracy of 71.70% on DepSeverity with BIN-4 suggests a favorable response to this prompt. Gemma 2 27b achieved accuracies of 62.40%, 65.40%, and 71.70% on Dreaddit, SDCNL Test, and DepSeverity, respectively.
Overall, most models showed a preference for the BIN-4 prompt, achieving the highest accuracy improvements, especially on the Dreaddit and DepSeverity datasets. This preference suggests that a structured and detailed prompt format generally enhances performance. The shift from BIN-1 to BIN-2 also led to significant improvements, highlighting the benefit of rephrasing tasks to focus on symptoms rather than direct diagnoses.
In terms of the overall best performing models, Gemma 7b as the top performer on the Dreaddit dataset with 76.70% accuracy (BIN-4), surpassing Mistral NeMo’s 74.50%. Other notable performances include Gemini 1.0 Pro, Mixtral 8x7b, and Mixtral 8x22b, which achieved accuracies of 72.20%, 69.50%, and 70.10% on BIN-4 and BIN-2, respectively. For the SDCNL dataset, Llama 3 70b matched GPT-4o’s 71.20% with BIN-2, while Gemma 7b reached 66.10% with prompt BIN-2. On the DepSeverity dataset, Mistral Medium achieved an accuracy of 75.10% (BIN-4), surpassing Llama 2 70b’s 73.58% (previous best on BIN-1). Several models, including Phi-3-small, Mixtral 8x7b, Mixtral 8x22b, Llama 3 70b, Gemma 2b, and Gemma 7b, approached or exceeded the 70% accuracy threshold, with Qwen 2 72b achieving the best performance at 76.60% (BIN-4).
Table 8: Merged Results of Prompt BIN-2/BIN-3/BIN-4 on Task 1
Model Dreaddit Test SDCNL Test DepSeverity
Claude 3 Haiku 56.60 / +5.20 / -6.60 / -5.60 / 61.802 60.70 / -0.80 / -6.90 / -6.40 / 60.701 69.20 / -0.10 / -15.80 / -15.20 / 69.201
Gemma 2b 49.49 / -3.09 / +10.21 / +0.51 / 59.703 5.0 / +2.00 / +39.00 / +45.00 / 50.004 13.0 / -4.00 / +51.00 / +39.00 / 64.003
Gemma 7b 54.12 / +13.68 / +11.88 / +22.58 / 76.704 54.21 / -5.31 / +11.89 / +0.89 / 66.103 57.10 / +11.88 / +11.89 / +6.20 / 68.993
Gemma 2 9b 56.8 / +1.30 / +3.80 / +4.20 / 61.004 65.2 / -6.80 / -5.70 / -2.20 / 65.201 68.5 / -0.40 / +1.80 / +3.20 / 71.704
Gemma 2 27b 55.6 / +3.00 / +4.10 / +6.80 / 62.404 63.5 / -1.00 / -1.30 / +1.90 / 65.404 66.0 / +4.10 / +4.10 / +5.70 / 71.704
Gemini 1.0 Pro 64.2 / +5.70 / +1.10 / +8.00 / 72.204 66.1 / -5.50 / -0.40 / -2.20 / 66.101 69 / +3.00 / +3.00 / +4.00 / 73.004
GPT-3.5-Turbo 71.4 / -5.50 / -18.50 / -17.10 / 71.401 66.2 / -2.00 / -14.30 / -13.20 / 66.201 71.4 / +0.10 / -14.60 / -10.00 / 71.502
Llama 3 8b 61.48 / +0.12 / -10.18 / -4.68 / 61.602 64.82 / -17.22 / -16.42 / -10.52 / 64.821 66.50 / +1.30 / -7.50 / -5.20 / 67.802
Llama 3.1 8b 53.30 / +4.00 / +3.70 / +10.80 / 64.104 48.50 / -2.20 / -26.80 / -5.00 / 48.501 63.30 / +5.30 / -0.30 / +6.80 / 70.104
Llama 2 13b 56.19 / +5.11 / -42.89 / +9.41 / 65.604 44.58 / +11.02 / -38.08 / -2.68 / 55.602 47.10 / +20.30 / -26.20 / +13.80 / 67.402
Llama 2 70b 72.70 / -11.70 / -21.50 / -22.60 / 72.701 59.5 / -2.45 / -7.55 / -8.05 / 59.501 73.58 / -6.58 / -21.08 / -20.28 / 73.581
Llama 3 70b 70.49 / -5.29 / -6.79 / +0.21 / 70.704 67.58 / +3.62 / -0.68 / -2.48 / 71.202 70.5 / +2.90 / +1.80 / +3.10 / 73.604
Llama 3.1 70b 68.30 / +4.30 / -5.70 / +0.70 / 72.602 67.10 / +1.10 / -5.90 / -1.80 / 68.202 69.80 / +0.60 / +1.00 / +2.20 / 72.004
Llama 3.1 405b 61.10 / +0.00 / -2.30 / +0.20 / 61.304 40.80 / -14.70 / -38.60 / -29.80 / 40.801 67.10 / +0.80 / -42.50 / -13.70 / 67.902
Mistral 7b 52.98 / -0.08 / +0.22 / +0.22 / 53.204 56.71 / +0.79 / -6.71 / -2.71 / 57.502 54.40 / +0.30 / -6.50 / +2.90 / 57.304
Mistral NeMo 74.50 / +0.20 / -23.50 / -23.20 / 74.702 58.20 / +5.90 / +2.90 / -3.30 / 64.102 64.30 / -0.50 / -7.90 / -6.20 / 64.301
Mixtral 8x7b 58.75 / +2.05 / +0.75 / +10.75 / 69.504 62.47 / -2.77 / +0.93 / +0.73 / 63.403 65.30 / -0.60 / +4.70 / +5.50 / 70.804
Mixtral 8x22b 57.81 / +12.29 / -3.81 / +2.19 / 70.102 62.59 / +3.51 / +0.31 / +1.41 / 66.102 51 / +13.00 / +14.00 / +21.00 / 72.004
Mistral Medium 57.5 / +3.50 / +6.50 / +9.40 / 66.904 67.6 / -0.90 / +1.50 / -1.50 / 69.102 59.7 / +9.30 / +13.60 / +15.40 / 75.104
Mistral Large 2 54.20 / +14.80 / -1.30 / +10.40 / 69.002 67.70 / +0.90 / -1.10 / -1.70 / 68.602 71.70 / -5.60 / +1.20 / +2.50 / 74.204
Phi-3-mini 58.9 / +4.80 / -1.80 / +0.80 / 63.702 59.3 / -1.70 / +1.00 / -6.50 / 60.303 70.5 / -2.10 / -4.00 / -6.40 / 70.501
Phi-3-small 55.1 / +2.30 / +4.50 / +5.50 / 60.604 65.4 / -10.70 / -0.30 / -6.10 / 65.401 63.4 / +2.90 / +9.20 / +7.90 / 72.604
Phi-3-medium 60.3 / -2.20 / -2.90 / -2.30 / 60.301 62.9 / -3.80 / +2.30 / -7.40 / 65.203 71.6 / -1.00 / -1.20 / +0.20 / 71.804
Qwen 2 72b 62.40 / +5.70 / -3.90 / +4.10 / 68.102 63.10 / -1.40 / -5.50 / -3.00 / 63.101 72.80 / +1.60 / -0.10 / +3.80 / 76.604
Notes: Results are formatted as: BIN-1 result / Change from BIN-2 / Change from BIN-3 / Change from BIN-4 / Highest Value across all prompts. The highest value across all models is underlined. The final best value for each model is provided with a superscript 1, 2, 3, or 4 indicating the corresponding prompt number from which the best result was obtained.
3.4 Disorder Severity Evaluation (task 2)
3.4.1 Results of prompt SEV-4 on the ZS Severity Evaluation task
An analysis of Table 9 reveals that the performance of each model on the severity task.
The GPT family of models exhibits strong performance across multiple tasks, with GPT-4-Turbo leading the pack. This model achieves the highest BA of 39.6% in the DepSeverity task and an outstanding 59.7% in DEPTWEET, making it the top performer for these datasets. Its MAE scores are notably low, with the minimum being 0.455 in DEPTWEET. The average BA and MAE for GPT-4-Turbo stand at 41.4% and 1.047, respectively, highlighting its robustness and reliability. GPT-4 also performs admirably, with a notable BA of 40.9% in the RED SAM Test and 29.0% in SAD, securing the second-best score in the latter. Consistent MAE scores, such as 0.827 in DepSeverity and 0.751 in RED SAM, contribute to an average BA of 40.48% and MAE of 1.141, affirming its strong performance across various datasets. While GPT-4o doesn’t lead in any specific dataset, it demonstrates competitive results with a BA of 35.6% in DepSeverity and 50.1% in DEPTWEET, complemented by a low MAE of 0.825. The GPT-3.5-Turbo model delivers moderate performance, peaking at 43.2% BA in DEPTWEET and maintaining consistent MAE scores, including 0.888 in DepSeverity. With an average BA of 31.28% and MAE of 1.120, it aligns closely with GPT-4o, though it doesn’t reach the top-tier performance seen in the latest models. The GPT-4o mini, a recent addition to the series, falls slightly behind the larger GPT-4o model, achieving an average BA of 34.10% and MAE of 1.230, indicating a slight gap in performance despite its more compact architecture.
The Claude models present a diverse range of capabilities. Claude 3 Sonnet stands out, particularly in the SAD dataset, where it achieves the highest BA of 29.7% and the lowest MAE of 1.404. It also performs strongly in DEPTWEET with a BA of 45.3% and an MAE of 0.683, resulting in an overall average BA of 35.73% and MAE of 0.975. This model demonstrates versatility and consistency across different datasets. Claude 3 Opus matches the highest BA in SAD (29.7%) and excels with a BA of 50.6% in DEPTWEET, along with a low MAE of 0.547, further emphasizing its effective performance. In contrast, Claude 3 Haiku shows generally poor results, with low average BA and high MAE, positioning it among the weaker models in this analysis. Claude 3.5 Sonnet, while consistent in some areas, does not match the performance of Claude 3 Opus or Sonnet in critical metrics, indicating a need for further refinement.
The Gemini and Gemma models offer notable performances, with Gemini 1.0 Pro achieving significant scores, including a BA of 54.2% in DEPTWEET and a second-best MAE of 0.729 in the RED SAM Test. Its average scores of 36.63% BA and 1.122 MAE reflect a strong overall showing. The Gemma models, particularly Gemma 2 9b, maintain consistent and competitive performance, with an average BA of 33.80% and MAE of 1.243, demonstrating their reliability across datasets. Gemma 2 27b also exhibits steady results, achieving an average BA of 29.98% and MAE of 1.234, making these models valuable for various applications.
Mistral models present a mixed bag of results, with Mistral Medium standing out for its BA of 39.0% in DepSeverity and the best MAE of 0.779. It maintains competitive average scores of 35.45% BA and 1.206 MAE, highlighting its consistent performance. Mistral Large and Mistral 7b deliver moderate performances, with BA scores of 31.5% and 31.9% in DepSeverity, respectively. Mistral NeMo, despite its smaller size, impresses with the best MAE of 0.639 in the RED SAM Test and a commendable MAE of 1.738 in SAD, resulting in an overall MAE of 1.006, which ranks among the best in its class. However, Mistral Large 2 underperforms compared to its predecessors, suggesting a need for optimization.
The Llama models display a broad spectrum of performance levels. Llama 3 70b emerges as the best among them, achieving a BA of 33.4% in DepSeverity and showing competitive scores across other datasets. In contrast, Llama 2 70b underperforms significantly, with a BA of just 12.7% in DepSeverity and high MAE scores, indicating challenges in handling the tasks. The newer Llama 3.1 family fails to surpass their Llama 3 counterparts, with the Llama 3.1 405b model performing similarly to the Llama 3.1 70b model, both of which do not meet the performance of the older Llama 3 70b model, reflecting a gap in expected advancements.
The Phi-3 models show varied outcomes. Phi-3-mini demonstrates moderate capabilities with a BA of 31.6% in DepSeverity and reasonable MAE scores. Phi-3-small, however, exhibits lower performance, with a BA of 25.5% in the same task. Phi-3-medium performs slightly better, achieving a BA of 32.1% in DepSeverity and maintaining competitive scores across other datasets.
Finally, the Qwen 2 72b model delivers moderate overall performance. While it doesn’t achieve standout results in any specific dataset, its performance is comparable to models like Llama 3 70b, Gemma 2 9b, or Phi-3-medium. It suggests a competent model that could benefit from further optimization to reach higher performance levels.
Table 9: Results of Prompt SEV-4 on Task 2 (ZS).
Model DepSeverity DEPTWEET RED SAM Test SAD Average
Claude 3 Haiku 27.6 / 1.006 34.3 / 0.878 36.0 / 0.798 11.4 / 2.781 27.3 / 1.366
Claude 3 Sonnet 32.8 / 0.960 45.3 / 0.683 35.1 / 0.853 29.7 / 1.404 35.7 / 0.975
Claude 3.5 Sonnet 30.0 / 0.970 39.0 / 0.767 34.8 / 0.868 15.4 / 2.474 29.8 / 1.270
Claude 3 Opus 31.1 / 0.873 50.6 / 0.547 37.3 / 0.788 29.7 / 1.776 37.2 / 0.996
Gemma 7b 30.2 / 0.987 34.2 / 0.923 33.2 / 0.905 15.4 / 3.287 28.3 / 1.525
Gemma 2 9b 33.2 / 0.846 50.4 / 0.559 34.8 / 0.775 16.8 / 2.793 33.8 / 1.243
Gemma 2 27b 30.6 / 0.907 39.8 / 0.728 33.6 / 0.843 15.9 / 2.460 30.0 / 1.234
Gemini 1.0 Pro 35.6 / 0.794 54.2 / 0.496 36.1 / 0.729 20.6 / 2.468 36.6 / 1.122
GPT-4o mini 35.3 / 0.858 45.7 / 0.649 36.5 / 0.813 18.9 / 2.559 34.1 / 1.230
GPT-3.5-Turbo 28.8 / 0.888 43.2 / 0.647 34.4 / 0.798 18.7 / 2.145 31.3 / 1.120
GPT-4o 35.6 / 0.825 50.1 / 0.577 38.5 / 0.762 18.2 / 2.671 35.6 / 1.209
GPT-4-Turbo 39.6 / 0.805 59.7 / 0.455 39.5 / 0.757 26.8 / 2.171 41.4 / 1.047
GPT-4 39.3 / 0.827 52.7 / 0.554 40.9 / 0.751 29.0 / 2.430 40.5 / 1.141
Llama 3 8b 26.1 / 0.929 37.0 / 0.769 31.4 / 0.769 21.2 / 1.655 28.9 / 1.281
Llama 3.1 8b 27.3 / 0.951 29.4 / 1.055 30.1 / 0.816 10.3 / 3.669 24.3 / 1.623
Llama 2 13b 26.3 / 1.045 32.8 / 0.924 30.6 / 0.966 8.6 / 3.932 24.6 / 1.717
Llama 2 70b 12.7 / 1.895 26.7 / 1.243 15.6 / 1.490 10.7 / 3.675 16.4 / 2.076
Llama 3 70b 33.4 / 0.819 42.3 / 0.674 33.5 / 0.832 16.1 / 2.648 31.3 / 1.243
Llama 3.1 70b 33.8 / 0.853 42.4 / 0.712 33.0 / 0.868 13.6 / 3.354 30.7 / 1.447
Llama 3.1 405b 33.0 / 0.877 41.8 / 0.812 31.6 / 0.900 13.6 / 2.890 30.0 / 1.370
Mistral 7b 31.9 / 1.055 44.4 / 0.767 37.6 / 0.797 14.6 / 2.811 32.1 / 1.358
Mistral NeMo 27.2 / 0.871 32.0 / 0.775 36.7 / 0.639 15.3 / 1.738 27.8 / 1.006
Mixtral 8x7b 35.0 / 0.939 44.7 / 0.789 32.8 / 0.871 20.2 / 2.180 33.2 / 1.195
Mixtral 8x22b 33.5 / 0.864 44.2 / 0.648 36.6 / 0.795 23.4 / 2.369 34.4 / 1.169
Mistral Medium 39.0 / 0.779 51.9 / 0.556 36.8 / 0.770 14.1 / 2.720 35.5 / 1.206
Mistral Large 31.5 / 0.898 43.2 / 0.680 35.2 / 0.807 16.8 / 2.566 31.7 / 1.238
Mistral Large 2 29.6 / 0.865 42.1 / 0.666 34.1 / 0.811 10.4 / 2.942 29.1 / 1.321
Phi-3-mini 31.6 / 1.015 40.7 / 0.865 35.1 / 0.865 13.6 / 3.011 30.3 / 1.439
Phi-3-small 25.5 / 0.897 31.9 / 0.786 33.6 / 0.757 9.6 / 2.691 25.2 / 1.282
Phi-3-medium 32.1 / 0.883 39.5 / 0.754 33.2 / 0.839 12 / 2.283 29.2 / 1.190
Qwen 2 72b 36.7 / 0.921 44.5 / 0.734 36.4 / 0.857 18.5 / 2.556 34.0 / 1.267
Notes: The results are written as follows: BA/ MAE. The best BA and MAE for each dataset are underlined separately. Each dataset may have different models underlined for BA and MAE.
3.4.2 Results of prompt SEV-4 on the FS Severity Evaluation task
An analysis of Table 10 reveals that most models generally improved on various datasets with the FS prompt. However, some models experienced minor decreases. Notably, even if a model improves across all datasets except SAD, the final average change might be negative due to SAD having the most significant impact, given its 10 severity levels.
Claude 3.5 Sonnet showed significant improvements, with an overall average improvement of +6.50% in BA and a reduction of -0.248 in MAE. In contrast, Claude 3 Sonnet experienced mixed results, with a +5.30% increase in BA and a -0.058 reduction in MAE in DEPTWEET, but a substantial decrease in SAD, resulting in an overall negative change.
Gemma models, particularly Gemma 2 27b, exhibited substantial positive changes, leading to an overall average improvement of +6.05% in BA and a reduction of -0.191 in MAE. Similarly, Gemma 2 9b showed moderate improvements with an overall +2.175% in BA and a reduction of -0.131 in MAE.
GPT models demonstrated positive trends, with GPT-4o showing consistent improvements across all metrics, resulting in an overall average improvement of +2.975% in BA and a reduction of -0.117 in MAE. GPT-4-Turbo had minor positive changes overall. GPT-4o mini experienced the greatest improvement of all the GPT models (+4.025% and -0.170).
Mistral models, particularly Mistral 7b, showed significant improvements with an overall average improvement of +6.52% in BA and a reduction of -0.345 in MAE. Mixtral models, such as Mixtral 8x7b, also performed well with an overall average improvement of +4.96% in BA and a reduction of -0.140 in MAE.
Some models experienced mixed or negative overall changes. Claude 3 Opus saw minor improvements but overall negative performance due to decreases in the RED SAM Test and SAD, leading to an overall average decrease of -2.70% in BA and an increase of +0.096 in MAE. Llama 2 13b suffered a major decrease in DepSeverity, resulting in an overall negative change.
Phi-3 models performed well with the FS prompt. Phi-3-mini showed significant improvements across all metrics, leading to an overall average improvement of +8.18% in BA and a reduction of -0.488 in MAE. Phi-3-medium also demonstrated consistent improvements.
Overall, utilizing the FS prompt with three examples generally enhanced performance across most models and datasets. While incorporating FS prompts can yield significant improvements, the degree of enhancement varies depending on the specific models and datasets. SAD, with its 10 severity levels, can disproportionately influence the final average, sometimes counteracting improvements observed in other datasets. Models like Claude 3.5 Sonnet, Gemma 2 27b, and Mixtral 8x7b demonstrated the most consistent and substantial enhancements. Notably, due to the inclusion of three additional example posts (in addition to the one provided for evaluation), the overall token count per evaluation is typically multiplied by 3-4.
With these enhancements, GPT-4-Turbo now excels in MAE across all datasets except RED SAM, where Gemini 1.0 Pro became the best. In terms of BA, the notable changes include Gemini 1.0 Pro’s dominance in RED SAM and Mistral 7b’s improved performance, making it the leading model for the SAD dataset.
Table 10: Results of Prompt SEV-4 on Task 2 (FS)
Model DepSeverity DEPTWEET RED SAM Test SAD Average
Claude 3 Haiku +0.30 / -0.059 -1.50 / -0.032 +0.10 / -0.080 -2.10 / +0.345 -0.80 / +0.043
Claude 3 Sonnet +1.80 / -0.042 +5.30 / -0.058 +0.40 / -0.055 -10.4 / +1.101 -0.73 / +0.237
Claude 3.5 Sonnet +1.50 / -0.089 +17.2 / -0.274 -0.40 / -0.053 +7.70 / -0.575 +6.50 / -0.248
Claude 3 Opus +0.90 / +0.025 +0.80 / -0.003 -3.40 / +0.031 -9.10 / +0.332 -2.70 / +0.096
Gemma 7b -1.00 / -0.060 +5.50 / -0.1530 -1.170 / +0.029 +6.33 / -0.596 +2.415 / -0.195
Gemma 2 9b +1.00 / -0.008 +5.80 / -0.068 +2.50 / -0.045 -0.60 / -0.419 +2.175 / -0.131
Gemma 2 27b +3.50 / -0.119 +12.5 / -0.196 +3.10 / -0.083 +5.10 / -0.366 +6.05 / -0.191
Gemini 1.0 Pro +2.70 / +0.007 -1.40 / +0.009 +7.4 / -0.086 +2.60 / -0.604 +2.825 / -0.168
GPT-4o mini +1.70 / -0.031 +7.10 / -0.113 +3.40 / -0.060 +3.90 / -0.475 +4.025 / -0.170
GPT-3.5-Turbo +4.20 / -0.001 -1.30 / +0.043 +3.00 / +0.003 +0.00 / +0.239 +1.475 / +0.071
GPT-4o +2.90 / -0.023 +5.50 / -0.047 +2.30 / -0.044 +1.20 / -0.354 +2.975 / -0.117
GPT-4-Turbo +3.60 / -0.022 +1.20 / -0.015 +3.50 / -0.083 -5.20 / -0.724 +0.78 / -0.211
GPT-4 +2.30 / -0.042 +8.50 / -0.100 +2.40 / -0.090 -10.5 / -0.244 +0.68 / -0.119
Mistral 7b +3.90 / -0.103 +7.97 / -0.131 +1.32 / -0.078 +12.9 / -1.066 +6.52 / -0.345
Mistral NeMo -1.50 / +0.045 +3.50 / -0.066 -0.60 / +0.007 +0.40 / +0.257 +0.45 / +0.061
Mixtral 8x7b -0.10 / -0.077 +8.00 / -0.169 +8.10 / -0.154 +3.83 / -0.159 +4.96 / -0.140
Mixtral 8x22b +2.10 / -0.035 +8.90 / -0.093 +2.92 / -0.045 -0.37 / -0.539 +3.39 / -0.178
Mistral Medium -1.80 / +0.049 -3.60 / +0.035 +5.60 / -0.095 +8.80 / -0.836 +2.25 / -0.212
Mistral Large +1.20 / -0.027 +7.70 / -0.126 +2.20 / -0.077 -4.10 / -0.094 +1.75 / -0.081
Mistral Large 2 123b +0.80 / -0.033 +6.00 / -0.105 +3.00 / -0.098 +2.90 / -0.621 +3.175 / -0.214
Llama 3 8b +0.10 / -0.046 -5.09 / +0.097 +1.90 / -0.096 -1.97 / -0.023 -1.27 / -0.017
Llama 3.1 8b +1.80 / -0.083 +3.50 / -0.173 +4.10 / -0.129 +3.00 / -0.876 +3.10 / -0.315
Llama 2 13b -24.4 / +1.379 +6.80 / -0.077 +0.16 / -0.049 +3.17 / +0.179 -3.57 / +0.358
Llama 2 70b -5.90 / +0.363 +4.00 / -0.113 -3.19 / +0.115 +9.61 / -1.274 +1.13 / -0.227
Llama 3 70b +1.10 / -0.023 +9.00 / -0.129 +3.22 / -0.089 +7.32 / -0.687 +5.16 / -0.232
Llama 3.1 70b +3.10 / -0.062 +10.00 / -0.182 +3.10 / -0.077 +4.80 / -0.808 +5.25 / -0.282
Llama 3.1 405b +2.00 / -0.078 +8.60 / -0.203 +4.10 / -0.150 +3.60 / -0.524 +4.58 / -0.239
Phi 3 Mini +6.80 / -0.202 +9.70 / -0.238 +5.40 / -0.213 +10.8 / -1.298 +8.18 / -0.488
Phi 3 Small +5.00 / -0.034 +8.00 / -0.122 +2.80 / -0.032 +3.90 / +0.253 +4.93 / +0.016
Phi 3 Medium +3.70 / -0.046 +6.50 / -0.107 +3.50 / -0.061 +6.70 / -0.128 +5.10 / -0.086
Qwen 2 72b +4.90 / -0.113 +7.70 / -0.158 +1.80 / -0.039 +7.80 / -0.347 +5.55 / -0.164
Notes: The results are written as follows: change to BA/ MAE.
3.5 Psychiatric Knowledge Assessment (task 3)
The previous experiments have highlighted each model’s effectiveness in classifying mental illnesses from social media posts. However, we still lack a fundamental understanding of each model’s inherent knowledge of psychiatry and various mental disorders. The results in figure 5 demonstrate the foundational psychiatric knowledge that each model possesses.
Llama 3.1 405b tops the performance chart, achieving an impressive accuracy of 91.2%, followed closely by Llama 3.1 70b with an accuracy of 89.50%. The smaller Llama 3.1 8b, however, lags behind with an accuracy of 68.0%. Despite this, it outperforms other models in its size category (under 14 billion parameters) such as Mistral 7b, Gemma 7b, and Phi-2, with a lead ranging from 8% to 20%. The older Llama 3 models all perform slightly worse compared to the more recently released models (e.g. Llama 3 70b actually beats GPT-4 even if only by 0.10%). Interestingly, it’s only in this task, which includes raw factual querying of the models, that the strength of the LLama 3.1 family and espcially the LLama 3.1 405b models stand out.
GPT-4o follows behind the LLama 3.1 models with an accuracy 88.6%, followed closely by GPT-4-Turbo at 87.8%, and GPT-4 at 85.5%. It’s clear that each successive iteration of the model, even while smaller and more efficient, is better trained and has a better overall understanding of psychiatric knowledge.
Claude 3.5 Sonnet follows, nearly matched by the Phi-3-medium model with accuracies of 82.5% and 82.2%, respectively. Both models are impressive for their size, but the Phi-3-medium model is especially notable, considering its relatively small size of 14 billion parameters, yet matching much larger models.
Claude 3 Opus slightly surpasses Mistral Large with an accuracy of 81.8% compared to 78.9%. The Gemma 2 27b follows closely with an accuracy of 78.6%, nearly matching Mistral Large despite its much smaller size.
Interestingly, the Mistral Medium model follows closely with an accuracy of 77.7%, matching the Mixtral 8x22b. Mistral Large 2 beats the previous Mistral Large and medium models but still loses out to the Claude 3 Opus model, and even to the GPT-4o mini model.
Phi-3-small, Gemma 2 9b, and Phi-3-mini follow in that order with accuracy scores of 77.3%, 76.0%, and 75.3%. Interestingly, Gemini 1.0 Pro’s accuracy matches that of Phi-3-mini at 75.3%, despite its much larger size.
Other models such as Mixtral 8x7b and GPT-3.5-Turbo show competitive performance, with accuracies of 75.3%, 73.2%, and 68.7%, respectively.
Claude 3 Haiku achieves an accuracy of 68.5%, while Claude 3 Sonnet performs worse with 62.7%. The Llama 2 models, particularly the 13b and 70b variants, struggle significantly in the knowledge task, achieving accuracies of 49.7% and 36.3%, respectively. Phi-2 also shows a subpar performance with an accuracy of 47.7%.
The general trend observed from these results is that while model size matters, more recent models perform significantly better than their older counterparts, even when there is a massive size differential (e.g., Phi-3-medium and Llama 3.1 70b, or the many other examples mentioned earlier). For models released around the same time, model size appears to be the deciding factor, with inconsistencies observed earlier (such as Gemma 2 9b outperforming Gemma 2 27b or Phi-3-mini outperforming Phi-2-small) being nearly absent in this task. It is likely that the more recent models were trained on larger and more diverse datasets, especially those including psychiatric data, thus enhancing their foundational knowledge in psychiatry.
Figure 5: Results of all models on Task 3.
Figure 6 further confirms this finding by illustrating the relationship between model accuracy and release date. The graph demonstrates a general upward trend in accuracy for models released over time, highlighting the ongoing advancements in model architecture, training data, and methodologies.
Notably, models like GPT-4 and GPT-4 Turbo, released in 2023, continue to perform exceptionally well despite their relative age. This success can be attributed to their substantial model sizes and the extensive datasets used for training. Interestingly, GPT-4o, although smaller than both GPT-4 and GPT-4 Turbo, outperforms them, likely due to enhancements in training techniques and optimization.
A similar trend is observed across other model families. The Llama 3.1 models show a marked improvement over the Llama 3 models, which themselves significantly outperform the older Llama 2 models. This progression underscores the importance of continued refinements in model development. Similarly, Mistral NeMo demonstrates a performance edge over earlier iterations, and Claude 3.5 Sonnet surpasses the previous Claude 3 models, reflecting the positive impact of recent advancements.
The graph also reinforces the observation that more recently released models generally outperform older, larger models. For instance, the Phi-3-medium model outperforms larger models such as Mistral Large and Mixtral 8x22b, suggesting that improvements in training quality and data diversity play a crucial role in model effectiveness.
However, within the same release period, model size remains a critical factor when training data quality is comparable. This is evident in the consistent performance gains seen with larger models, highlighting that, while newer training techniques and data can enhance smaller models, the additional capacity provided by larger models often leads to superior accuracy.
2022-11-282023-03-142023-07-182023-09-272023-11-142023-12-092023-12-122023-12-232024-02-212024-02-262024-03-142024-04-182024-05-132024-06-062024-06-202024-06-272024-07-022024-07-182024-07-232024-07-2430303030404040405050505060606060707070708080808090909090100100100100Claude 3 HaikuClaude 3 SonnetClaude 3.5 SonnetClaude 3 OpusGemma 2bGemma 7bGemma 2 9bGemma 2 27bGemini 1.0 ProGPT-4o miniGPT-3.5-TurboGPT-4oGPT-4-TurboGPT-4Llama 3 8bLlama 2 13bLlama 2 70bLlama 3 70bLlama 3.1 8bLlama 3.1 70bLlama 3.1 405bMistral 7bMistral NeMoMixtral 8x7bMixtral 8x22bMistral MediumMistral LargeMistral Large 2Phi-2Phi-3-miniPhi-3-smallPhi-3-mediumQwen 2 72bRelease DateAccuracyClaudeGemma/GeminiGPTLlamaMistralPhiQwen
Figure 6: Results of all models on Task 3, sorted by model release date
4 Additional Investigations
In this section, we discuss side experiments that we performed while on the course of this larger analysis that are interesting but not core to this study.
4.1 Repetition Line Experiment
In this experiment, we evaluate the effect of the repetition line presented in the SEV-4 prompt (Figure 1). The effect of the line is observed in a binary task in isolation of any other modifications. The line is appended to the end of the binary BIN-1 prompt as showcased in Table 11.
Table 11: Prompt BIN-1.1
Prompts BIN-1.1
As a psychologist, analyze the provided social media post(s).
Is the poster depressed?
(Answer with Yes or No only without explaining your reasoning)
(Please return only the label (Yes or No) and no other text)
The results in Table 12 reveal a consistent trend of marked improvements in nearly all models with the improvements ranging from 1% to an upmost of 14%. Interestingly, the models that had the largest improvements (>10%) are all Mistral-based, which suggests that this particular family of models, whether for training reasons or how they were initially trained, seem to prefer reiterating the instructions or the important information in a prompt. Other model families generally still had respectable improvements of around 1-2%, which suggests that even other models prefer reiterating certain parts of the prompt, with only Claude 3 Sonnet and GPT-3.5-Turbo having a decrease in performance.
Table 12: Effect of Repetition Line in Binary Prompt BIN-1 on DepSeverity (Binary - 1000 Samples)
Model BIN-1 BIN-1.1
Claude 3 Haiku 69.20% 69.10% (-0.10%)
Claude 3 Sonnet 69.10% 66.40% (-2.70%)
Gemma 7b 54.10% 55.00% (+0.90%)
Gemma 2 9b 68.50% 67.60% (-0.90%)
Gemma 2 27b 66.00% 68.50% (+2.50%)
Gemini 1.0 Pro 69.00% 71.40% (+2.40%)
GPT-3.5-Turbo 71.40% 67.00% (-4.40%)
Llama 3 8b 61.48% 66.80% (+5.32%)
Llama 3.1 8b 63.30% 66.20% (+2.9%)
Llama 2 70b 72.70% 75.00% (+2.30%)
Llama 3 70b 70.49% 71.90% (+1.41%)
Llama 3.1 70b 69.80% 69.90% (+0.1%)
Llama 3.1 405b 67.10% 68.70% (+1.6%)
Mistral 7b 52.98% 64.40% (+11.42%)
Mistral NeMo 64.30% 61.10% (-3.20%)
Mixtral 8x7b 58.75% 69.30% (+10.55%)
Mixtral 8x22b 57.81% 71.90% (+14.09%)
Mistral Medium 59.70% 71.10% (+11.40%)
Mistral Large 2 54.20% 72.50% (+18.30%)
Phi-3-mini 70.50% 70.10% (-0.40%)
Phi-3-small 63.40% 70.20% (+6.80%)
Phi-3-medium 71.60% 72.50% (+0.90%)
Qwen 2 72b 72.80% 75.10% (+2.3%)
4.2 Knowledge Reinforcement Experiment
In this experiment, we sought to determine whether providing models with information they had previously generated would enhance their performance. This idea stemmed from an earlier study (not included in this paper) where we provided models with the DSM-5 criteria for depression. Interestingly, some models improved in accuracy, while others experienced a decline. To investigate this phenomenon more systematically, we designed the Knowledge Reinforcement Experiment.
The experiment was conducted in two steps. First, each LLM was prompted to list symptoms it associates with depression. The prompt used for this step is presented in Table 13.
Table 13: Prompt used for extracting disorder symptoms from LLM s
Give me a detailed list of symptoms that indicate whether a person is depressed or not.
In the second step, the list of symptoms provided by each model was appended to a modified version of the BIN-1 prompt, as shown in Table 14. This modified prompt, which we call the Knowledge Reinforcement Prompt, included a repetition line designed to help the models adhere more closely to the instructions, particularly since the detailed list of symptoms significantly expanded the prompt. The longer prompts occasionally caused models (notably the Mistral Large and Claude 3 Opus) to disregard some instructions, such as providing a simple "Yes" or "No" answer without explanation.
Table 14: Prompt BIN-1 + Disorder symptoms extracted from the same LLM that is being prompted
Knowledge Reinforcement Prompt
As a psychologist, analyze the provided social media post(s).
Is the poster depressed?
(Answer with Yes or No only without explaining your reasoning)
(Please return only the label (Yes or No) and no other text)
List of depression symptoms:
Table 15 presents the results of this experiment. The main goal is to observe whether a model would benefit from being provided information it very clearly knows, given that it provided us that information to begin with.
The results lack a clear pattern, with some models showing a moderate increase in accuracy, notably GPT-3.5-Turbo and Claude 3 Sonnet. However, most smaller LLM s experienced a decrease in accuracy, some quite significantly. We are unsure why smaller LLM s were negatively affected, as the additional information should have been neutral or even beneficial. One hypothesis is that the lengthy prompts with extensive information might have distracted them from the main instructions and task, potentially leading to poorer performance.
Table 15: Results of the Knowledge Reinforcement experiment compared to the baseline BIN-1 prompt results
Model Prompt BIN-1.1 (baseline) Knowledge Reinforcement Prompt
Claude 3 Haiku 69.1% -9.7%
Claude 3 Sonnet 66.4% +4.2%
GPT-3.5-Turbo 67% +2.8%
Gemini 1.0 Pro 71.4% +0.2%
Mistral Medium 71.1% +1.3%
Mixtral 8x22b 71.9% -3.2%
Llama 3 70b 71.9% +0.9%
Llama 2 70b 75% -4.1%
Mixtral 8x7b 69.3% -4.3%
Llama 3 8b 66.8% -8.5%
Gemma 7b 55% 0%
Mistral 7b 64.4% -0.6%
5 Challenges, Conclusions and Future Directions
5.1 Challenges
While the potential of LLM s in the realm of mental health is promising, our investigation into their capabilities for binary disorder detection faced several challenges. These obstacles arose from both the limitations of available datasets and the inherent characteristics of the LLM s themselves. Understanding these challenges is crucial for interpreting the results of our evaluation and guiding future research in this domain.
5.1.1 Datasets Based Challenges
The ideal evaluation would involve clinical reports containing patient summaries and diagnoses, as these most accurately reflect an individual’s physical and mental health. However, due to the confidential nature of such data, access to these datasets is either extremely difficult or nearly impossible without a close working relationship with a hospital or mental health institution. As a result, we relied on publicly available social media datasets, which, while more accessible, are inherently less accurate.
These social media datasets presented quality concerns. Many labels were based on context (e.g., subreddit of origin), leading to potential noise in the data. For instance, posts from r/Depression may not always indicate depression, but could be supportive messages or questions. We prioritized datasets with human supervision to mitigate this issue.
However, even with human-labeled data, psychiatric evaluation remains a subjective process. Trained professionals may have differing opinions, which can introduce variability into the data. As a result, even datasets labeled by experts may still contain some degree of noise due to these differences in interpretation.
5.1.2 LLM Based Challenges
The use of LLM s presents several challenges. One significant issue is the cost associated with models like GPT-4 and Claude Opus, which limited our experiments to samples from test sets. This high expense raises concerns about the scalability of these models for real-world applications.
Another challenge is the context limit of some models, such as Phi-2, which has a 2048-token limit. This constraint made it difficult to handle datasets with very long posts. However, with newer models featuring larger context windows, this limitation is expected to become less of a concern in the future.
API filtering presents a significant challenge. For example, Gemini 1.0 Pro’s API has stringent filtering policies that block prompts mentioning sensitive topics such as pedophilia, sexual abuse, or grooming. This restriction limits the model’s applicability in mental health contexts, where discussing sensitive issues is often necessary. Consequently, this filtering issue was the primary reason other Gemini models, such as Gemini 1.5 Flash/Pro, were not tested in our experiments.
Ethical guards implemented in some models, such as the Llama 3.1, have also been problematic. These safeguards can significantly hinder performance in mental health contexts compared to their older Llama 3 counterparts.
Additionally, some models exhibit stubborn behavior, resisting the provision of explanations despite clear instructions. Mistral models, in particular, have shown a higher level of stubbornness compared to GPT models. This inconsistency in following instructions is concerning and may be linked to the training data or processes used.
Finally, the sensitivity of models to prompt changes has been another issue. Even minor modifications to prompts can lead to excessive sensitivity. For example, adding redundant information notably affected the performance of models like Llama 3 70b and Gemini 1.0 Pro, highlighting the importance of careful prompt design.
5.2 Future Directions
Building upon the insights gained from our investigation into the use of LLM s for mental health applications, several promising avenues for future research have emerged. These directions are aimed at enhancing the interpretability, performance, and applicability of LLM s in real-world mental health contexts, while also addressing some of the key challenges identified in our study.
One promising area of exploration is the application of CoT prompting to enhance the interpretability and decision-making capabilities of LLM s in psychiatric contexts. CoT prompting involves structuring model prompts in a way that encourages step-by-step reasoning, which could be particularly beneficial for understanding complex mental health issues. This approach could help clarify the rationale behind model outputs, making it easier to validate and trust the decisions made by LLM s. Investigating how CoT prompting affects the accuracy and reliability of mental health diagnostics will be crucial for deploying LLM s in real-world psychiatric applications.
Another innovative approach to improving LLM performance involves the use of Retrieval-Augmented Generation (RAG) techniques. By integrating LLM s with a database of relevant knowledge, such as books, case studies, or even anonymized medical records, RAG could provide additional context and grounding to the model’s responses. This method can enhance the model’s ability to generate more informed and contextually accurate outputs, thus improving diagnostic accuracy and relevance.
Fine-tuning LLM s on specific psychiatric datasets represents another important area for future investigation. Tailoring models to domain-specific data could potentially improve their performance and generalizability across different mental health conditions. Evaluating the impact of fine-tuning on diagnostic accuracy and model robustness will provide insights into how LLM s can be optimized for diverse psychiatric tasks and populations.
Expanding the evaluation of LLM s to include other psychiatric tasks, such as dialogue safety and empathy, is essential for a comprehensive understanding of their capabilities. Publicly available datasets for these tasks can provide valuable insights into how well LLM s perform beyond binary disorder detection. Additionally, examining how LLM s handle the diagnosis of multiple co-occurring mental disorders could provide a more nuanced view of their diagnostic capabilities and limitations.
Furthermore, future research should include evaluating LLM s with datasets in other languages, as well as those involving different modalities such as audio, images, and videos. With advancements in multimodal models like GPT-4o and Llama 3.1, integrating and analyzing data from diverse sources can enhance diagnostic accuracy and offer a richer understanding of mental health conditions across various contexts.
By pursuing these directions, researchers can address current challenges and push the boundaries of LLM applications in mental health, ultimately leading to more effective and ethically sound tools for psychiatric care.
5.3 Conclusions
In this study, we conducted an extensive evaluation of the capabilities of LLM s in the context of mental health tasks, specifically focusing on the detection and evaluation of mental disorders from social media posts. Our investigation encompassed a wide range of LLM s, including state-of-the-art models like GPT-4 and Claude 3.5, as well as smaller, more accessible models such as Gemma 2, Phi-3, LLama 3/3.1 and others. We employed diverse prompt templates and evaluation metrics to assess their performance across three primary tasks: binary disorder detection, disorder severity evaluation, and psychiatric knowledge assessment.
Our findings revealed that LLM s hold significant promise in assisting with mental health tasks. In the binary disorder detection task, models like GPT-4 and Llama 3 exhibited superior performance, achieving accuracies of up to 85% on certain datasets. The choice of prompt template played a crucial role, with some models favoring more structured prompts while others performed better with open-ended ones. The disorder severity evaluation task highlighted the benefits of few-shot learning, where providing models with a few examples before making predictions led to improved performance in most cases. In the psychiatric knowledge assessment task, recent models generally outperformed older, larger counterparts, demonstrating the advancements in model architecture and training data.
However, the study also identified challenges that must be addressed to effectively deploy LLM s in mental health applications. The limitations of social media datasets, such as potential noise and subjectivity in labeling, and the high costs associated with large models, pose significant concerns. Additionally, ethical restrictions imposed by LLM providers on sensitive queries complicate accurate evaluations in mental health contexts. Despite these hurdles, the potential of LLM s in mental healthcare is undeniable. Future research should focus on improving interpretability and decision-making through techniques like chain-of-thought prompting and retrieval-augmented generation. Additionally, fine-tuning models on specific psychiatric datasets and exploring multimodal data integration can further enhance their utility in this critical domain.
6 Acknowledgement
We gratefully acknowledge and thank the authors of [29] for providing the DEPTWEET dataset.